### v1

# Deblur4DGS: 4D Gaussian Splatting from Blurry Monocular Video

Abstract
Recent 4D reconstruction methods have yielded impressive results but rely on sharp videos as supervision. However, motion blur often occurs in videos due to camera shake and object movement, while existing methods render blurry results when using such videos for reconstructing 4D models. Although a few NeRF-based approaches attempted to address the problem, they struggled to produce high-quality results, due to the inaccuracy in estimating continuous dynamic representations within the exposure time. Encouraged by recent works in 3D motion trajectory modeling using 3D Gaussian Splatting (3DGS), we suggest taking 3DGS as the scene representation manner, and propose the first 4D Gaussian Splatting framework to reconstruct a high-quality 4D model from blurry monocular video, named Deblur4DGS. Specifically, we transform continuous dynamic representations estimation within an exposure time into the exposure time estimation. Moreover, we introduce exposure regularization to avoid trivial solutions, as well as multi-frame and multi-resolution consistency ones to alleviate artifacts. Furthermore, to better represent objects with large motion, we suggest blur-aware variable canonical Gaussians. Beyond novel-view synthesis, Deblur4DGS can be applied to improve blurry video from multiple perspectives, including deblurring, frame interpolation, and video stabilization. Extensive experiments on the above four tasks show that Deblur4DGS outperforms state-of-the-art 4D reconstruction methods. The codes are available at https://github.com/ZcsrenlongZ/Deblur4DGS.
1 Introduction
Figure 1: Performance and rendering speed comparisons for novel-view synthesis on the 720 $\times$ 1080 images. Deblur4DGS produces better results with real-time rendering speed.
Substantial efforts have been made for 4D reconstruction, which has extensive applications in augmented reality and virtual reality. To model static scenes, Neural Radiance Field (NeRF) [52] and 3D Gaussian Splatting (3DGS) [27] propose implicit neural representation manner and explicit Gaussian ellipsoids one, respectively. To model dynamic objects, implicit neural fields [48, 108, 90, 84, 5, 60, 87] and explicit deformation [10, 8, 49, 89, 25, 30, 44, 40, 79] are suggested for motion representation. While achieving great progress, most methods rely on synchronized multi-view videos. They yield unsatisfactory results when applied to monocular video, where dynamic objects are only observed once at each timestamp. To alleviate the under-constrained nature of the problem, recent studies have introduced data-driven priors, such as depth maps [35, 88], optical flows [16, 76, 17, 19, 46, 108, 35, 80], tracks [79, 65, 70, 36], and generative models [86, 8, 91, 93] to improve 4D reconstruction.
Unfortunately, motion blur often arises in the videos due to camera shake and object movement, while the above methods heavily rely on sharp videos for supervision. When reconstructing the 4D scene from the blurry video, these methods can render blurry results. The first step to solving this problem is to deal with camera motion blur, which is relatively simple. Some NeRF-based [32, 50, 78, 33, 39] and 3DGS-based [102, 58, 7, 55] methods have suggested jointly optimizing 3D static scene representation and camera poses within the exposure time by calculating the reconstruction loss between the synthetic blurry images and the input blurry frames. In contrast, the object motion blur is more challenging to address, as the solution has to estimate continuous and sharp dynamic representations within the exposure time to simulate blurry frames. As far as we know, only DybluRF [72, 4] attempted this with NeRF but struggled to produce high-quality and real-time rendering results, as shown in Fig. 1.
In this work, we propose taking 3DGS [27] as the scene representation manner to explore the problem, driven by two main motivations. First, the real-time rendering capability of 3DGS and its successful application in 4D reconstruction make this method highly promising. Second, the explicit representation property of 3DGS makes it easier to model 3D object motion than NeRF. In particular, recent advances [30, 79, 65, 70] in reconstructing 3D motion trajectories present an opportunity to simplify this continuous dynamic representations estimation problem. Based on this, the challenging problem can be transformed into simple exposure time estimation, without the extra motion trajectory modeling in DybluRF [72, 4]. With the estimated exposure time, continuous dynamic representations can be obtained by directly interpolating between representations at the nearest integer timestamps.
Specifically, we name it Deblur4DGS, which is the first Gaussian Splatting framework designed for 4D reconstruction from blurry monocular video. For the static scene, we present a camera motion predictor to learn camera poses at exposure start and end, which is jointly optimized with static Gaussians. For the dynamic objects, we optimize learnable exposure time parameters and dynamic Gaussians of the integer timestamps, simultaneously. Then continuous camera poses and dynamic Gaussians within exposure time can be obtained by interpolation, and they are used to render continuous sharp frames to calculate the reconstruction loss. Moreover, the exposure regularization term is introduced to avoid trivial solutions, while multi-frame and multi-resolution consistency regularization terms are used to alleviate artifacts. Furthermore, existing 4D reconstruction methods generally select Gaussians at a single timestamp as canonical Gaussians (e.g., the ones at the first timestamp). However, it may produce results with missing details in scenes with large motion, especially when processing blurry videos with a low frame rate. To alleviate this issue, we suggest variable canonical Gaussians as time progresses based on the image blur level. Gaussians corresponding to the sharper frame are selected as the canonical ones for better blur removal, and each canonical Gaussian is only used for some nearby timestamps to reduce the difficulty of modeling large motion.
Blurry videos suffer from not only motion blur, but also low frame rates and scene shake generally. Beyond novel-view synthesis, the optimized Deblur4DGS can also be applied to address these problems, achieving deblurring, frame interpolation, and video stabilization. We evaluate the 4D reconstruction quality of Deblur4DGS from all four perspectives. Extensive experiments demonstrate that Deblur4DGS outperforms state-of-the-art 4D reconstruction methods quantitatively and qualitatively while maintaining real-time rendering speed. Furthermore, Deblur4DGS has competitive capabilities in comparison with task-specific video processing models trained in a supervised manner.
The main contributions can be summarized as follows:
  • We propose Deblur4DGS, the first 4D Gaussian Splatting framework specially designed to reconstruct a high-quality 4D model from blurry monocular video.
  • We propose transforming dynamic representation estimation into exposure time estimation, where a series of regularizations are suggested to tackle the under-constrained optimization and blur-aware variable canonical Gaussians is present to better represent objects with large motion.
  • Extensive experiments on novel-view synthesis, deblurring, frame interpolation, and video stabilization tasks show that Deblur4DGS significantly outperforms state-of-the-art 4D reconstruction methods.
2 Related Work
2.1 Image and Video Deblurring
Deep learning-based image [73, 62, 38, 29, 82, 92, 37, 99, 43, 101] and video [23, 54, 69, 83, 106, 42, 41, 56, 95, 66, 105, 24, 104, 6, 96] deblurring methods have been widely explored. Compared to image deblurring methods, video ones can leverage temporal clues between consecutive frames for more effective restoration. For example, DBLRNet [96] employs spatial-temporal 3D convolution for spatial and temporal features aggregation. DSTNet [56] develops a deep discriminative spatial and temporal network. BasicVSR++ [6] improves feature fusion with second-order feature propagation and flow-guided deformable alignment. BSSTNet [95] introduces a blur map to sufficiently utilize the entire video, achieving recent state-of-the-art. When reconstructing from a blurry video, pre-processing it with the 2D deblurring method is a straightforward manner. However, 2D deblurring methods cannot perceive 3D structures and maintain scene geometric consistency, leading to unsatisfactory scene reconstruction.
Figure 2: (a) Training of Deblur4DGS. When processing $t$-th frame, we first discretize its exposure time into $N$ timestamps. Then, we estimate continuous camera poses $\{\mathbf{P}_{t,i}\}_{i=1}^{N}$ (see Sec. 4.1) and dynamic Gaussians $\{\mathbf{D}_{t,i}\}_{i=1}^{N}$ (see Sec. 4.2) within exposure time. Next, we render each latent sharp image $\hat{\mathbf{I}}_{t,i}$ with the camera pose $\mathbf{P}_{t,i}$, dynamic Gaussians $\mathbf{D}_{t,i}$ and static Gaussians $\mathbf{S}$. Finally, $\{\hat{\mathbf{I}}_{t,i}\}_{i=1}^{N}$ are averaged to obtain the synthetic blurry image $\hat{\mathbf{B}}_{t}$, which is used to calculate the reconstruction loss $\mathcal{L}_{rec}$ with the given blurry frame $\mathbf{B}_{t}$. To regularize the under-constrained optimization, we introduce exposure regularization $\mathcal{L}_{e}$, multi-frame consistency regularization $\mathcal{L}_{mfc}$ and multi-resolution consistency regularization $\mathcal{L}_{mrc}$ (see Sec. 4.3). (b) Rendering of Deblur4DGS. Deblur4DGS produces the sharp image with user-provided timestamp $t$ and camera pose $\mathbf{P}_{t}$.
2.2 3D and 4D Reconstruction
To reconstruct 3D models, NeRF [52] and 3DGS [27] introduce implicit neural representation manner and explicit Gaussian ellipsoids one respectively, where the latter generally achieves better results in both image quality and rendering speed. To reconstruct 4D models, most works [60, 87, 5, 15, 13, 45, 18, 103, 14, 68, 11, 48, 49, 89, 30, 20, 44, 40, 67, 74, 84, 90, 1, 51, 10, 8, 25] incorporate implicit neural fields and explicit deformation for motion representation. For example, DeformableGS [90] achieves it with multi-layer perceptrons (MLPs). 4DGasusians [1] improves it with a multi-resolution HexPlane [5]. E-D3DGS [84] introduces per-Gaussian embeddings and temporal embeddings. Moreover, to better reconstruct from monocular video, some studies enhance 4D reconstruction with data-driven priors, such as depth maps [35, 88], optical flows [16, 76, 17, 19, 46, 108, 35, 80], tracks [79, 65, 70, 36], and generative models [86, 8, 91, 93, 98].
Note that these methods heavily rely on high-quality sharp videos for supervision and perform poorly when facing blurry inputs. To process camera motion in static areas, recent works [32, 50, 78, 33, 39, 31, 102, 58, 34, 7, 55] suggest jointly optimizing the scene representation and recovering the camera poses within the exposure time. To process object motion blur in dynamic scenes, DybluRF [72, 4] incorporates object motion blur formation into dynamic model optimization but faces challenges in producing high-quality images and achieving real-time rendering. In this work, with 3DGS [27] as the scene representation manner, we develop Deblur4DGS to reconstruct a high-quality 4D model from a blurry video.
3 Preliminary
3.1 4D Gaussian Splatting
A 3D Gaussian [27] is parameterized by $\{\mathbf{x},\mathbf{r},\mathbf{s},\mathbf{o},\mathbf{c}\}$, where $\mathbf{x}$ characterizes the center position in the world space, rotation matrix $\mathbf{r}$ and scale matrix $\mathbf{s}$ define the shape, $\alpha$ is opacity, and spherical harmonics (SH) coefficients $\mathbf{c}$ represent the view-dependent color.
In 4D Gaussian Splatting (4DGS), the static and dynamic regions are usually processed, respectively. Static regions can be represented by a set of 3D Gaussians, named $\mathbf{S}$. For the dynamic areas, 4DGS generally selects a timestamp (e.g., the first timestamp) and represents the objects by canonical dynamic Gaussians, i.e., $\mathbf{C}$. Then, $\mathbf{C}$ is deformed to other timestamps by an implicit neural field [48, 108, 90, 84, 5, 60, 87] or explicit manner [10, 8, 49, 89, 25, 30, 44, 40, 79] for motion representation. Denote by $\mathbf{D}_{t}$ the dynamic Gaussians at $t$-th timestamp, it can be written as,
$\mathbf{D}_{t}=\mathcal{F}(\mathbf{C},t;\Theta_{\mathcal{F}}).\vspace{-2mm}$ (1)
$\mathcal{F}$ is the deformation operation with parameters $\Theta_{\mathcal{F}}$. The Gaussians for $t$-th timestamp is the union of $\mathbf{S}$ and $\mathbf{D}_{t}$.
Collectively, 4DGS models a scene with static Gaussians $\mathbf{S}$, canonical dynamic Gaussians $\mathbf{C}$, and a deformation operation $\mathcal{F}$. With the provided camera pose, the Gaussians at $t$-th timestamp $\mathbf{D}_{t}$ can be projected into 2D spaces and rasterized to obtain the corresponding image.
3.2 Motion Blur Formation
Motion blur occurs due to camera shake and object movement, which can be simply regarded as the integration of a latent sharp image sequence [53], i.e.,
[…]
$\mathbf{B}(u,v)\in\mathbb{R}^{H\times W\times 3}$ is the blurry image and $\mathbf{I}_{t}(u,v)$ is the latent sharp one at $t$-th timestamp. $(u,v)$ is pixel location, $\tau$ is the camera exposure time, and $\phi$ is a normalization factor. To approximate the integral operation, recent works [102, 72, 78] divide the exposure time into $N$ timestamps and regard the blurry image as the average of $N$ sharp images, i.e.,
$\mathbf{B}(u,v)\approx\frac{1}{N}\sum_{i=0}^{N-1}\mathbf{I}_{i}(u,v).\vspace{-2mm}$ (3)
In this work, we reconstruct a high-quality 4D model from a blurry video by incorporating the physical motion blur formation into model optimization.
4 Proposed Method
Let $\{\mathbf{B}_{t}\}_{t=1}^{T}$ and $\{\mathbf{M}_{t}\}_{t=1}^{T}$ denote a blurry video with $T$ timestamps and the corresponding masks indicating dynamic areas, respectively. As shown in Fig. 2(a), when processing $t$-th frame, we first evenly divide its camera exposure time into $N$ timestamps. Then, we estimate continuous camera poses $\{\mathbf{P}_{t,i}\}_{i=1}^{N}$ and dynamic Gaussians $\{\mathbf{D}_{t,i}\}_{i=1}^{N}$ to simulate camera shake and object movement. Next, we render each sharp image $\hat{\mathbf{I}}_{t,i}$ with the corresponding camera pose $\mathbf{P}_{t,i}$, dynamic Gaussians $\mathbf{D}_{t,i}$ and static Gaussians $\mathbf{S}$. After that, we average $\{\hat{\mathbf{I}}_{t,i}\}_{i=1}^{N}$ to obtain the synthetic blurry image $\hat{\mathbf{B}}_{t}$, which is used to calculate the reconstruction loss $\mathcal{L}_{rec}$ with the given blurry frame $\mathbf{B}_{t}$, i.e.,
$\mathcal{L}_{rec}=(1-\beta)\mathcal{L}_{1}(\hat{\mathbf{B}}_{t},\mathbf{B}_{t})+\beta\mathcal{L}_{ssim}(\hat{\mathbf{B}}_{t},\mathbf{B}_{t}).\vspace{-1mm}$ (4)
$\mathcal{L}_{1}$ and $\mathcal{L}_{ssim}$ are $\ell_{1}$ loss and SSIM [81] loss, respectively. $\beta$ is set to $0.2$. The settings all follow 3DGS [27].
4.1 Continuous Camera Poses Estimation
To estimate continuous camera poses, recent methods [102, 58, 7, 55] directly optimize exposure start and end poses (i.e., $\mathbf{P}_{t,1}$ and $\mathbf{P}_{t,N}$). Then, the linear interpolation is performed between $\mathbf{P}_{t,1}$ and $\mathbf{P}_{t,N}$ in the Lie-algebra of $\mathbf{SE}(3)$ to obtain the camera pose at $i$-th intermediate timestamp (i.e., $\mathbf{P}_{t,i}$). It can be written as,
$\mathbf{P}_{t,i}=\mathbf{P}_{t,1}\odot\texttt{exp}(\frac{i-1}{N-1}\odot\texttt{log}(\frac{\mathbf{P}_{t,N}}{\mathbf{P}_{t,1}})).\vspace{-2mm}$ (5)
exp and log are exponential and logarithmic functions, respectively. $\odot$ is a pixel-wise multiply operation.
We follow the manner but deploy a tiny MLP as the camera motion predictor (see details in Suppl.) for more stable optimization. We pre-train the camera motion predictor and static Gaussians $\mathbf{S}$ with static reconstruction loss $\mathcal{L}_{rec}^{s}$, i.e.,
$\mathcal{L}_{rec}^{s}=(1-\beta)\mathcal{L}_{1}(\hat{\mathbf{B}}_{t}^{s},\mathbf{B}_{t}^{s})+\beta\mathcal{L}_{ssim}(\hat{\mathbf{B}}_{t}^{s},\mathbf{B}_{t}^{s}).\vspace{-1mm}$ (6)
$\hat{\mathbf{B}}_{t}^{s}=(1-\mathbf{M}_{t})\odot\hat{\mathbf{B}}_{t}$ and $\mathbf{B}_{t}^{s}=(1-\mathbf{M}_{t})\odot\mathbf{B}_{t}$ are the static areas of $\hat{\mathbf{B}}_{t}$ and $\mathbf{B}_{t}$, respectively.
4.2 Continuous Dynamic Gaussians Estimation
We first introduce blur-aware variable canonical Gaussians for better dynamic representation at integer timestamps. Then, we describe Gaussian deformation manner. Finally, we detail how to take learnable exposure time parameters to obtain continuous dynamic Gaussians within exposure time.
Blur-Aware Variable Canonical Gaussians. Existing 4D reconstruction methods generally select a single canonical Gaussians $\mathbf{C}$ across the entire video, which may produce results with missing details in scenes with large motion. To alleviate the issue, we suggest variable the canonical Gaussians as time progresses. In such case, the $k$-th canonical Gaussians $\mathbf{C}_{k}$ is only used for some nearby timestamps thus reducing the difficulty of modeling large motion. One way to achieve this is to uniformly divide the video into $L$ segments and select $\mathbf{C}_{k}$ for $k$-th segment. Although it improves performance, selecting the one corresponding to the sharper frame is better for blur removal. In particular, we first uniformly divide the video into $L$ segments and calculate the blur level $b_{t}$ of dynamic areas for $t$-th frame following  [2, 63]. It can be written as,
$b_{t}=\sum_{(u,v)\in\mathbf{M}_{t}}(\Delta\mathbf{B}_{t}(u,v)-\overline{\Delta\mathbf{B}_{t}})^{2},\vspace{-2mm}$ (7)
where $\mathbf{M}_{t}$ indicates dynamic areas. $\Delta\mathbf{B}_{t}$ is the image Laplacian and $\overline{\Delta\mathbf{B}_{t}}$ is its mean value. The larger $b_{t}$ is, the sharper the frame is. To make the start and end frame of the segment as sharp as possible, we look for the sharp frame among their surrounding $H$ frames and redefine them as the start and end of the current segment. After that, we select the Gaussians corresponding to the sharpest frame in each segment as its canonical ones. Gaussian Deformation. We follow Shape-of-Motion [79] to deform dynamic Gaussians, where a set of rigid transformation matrices are deployed. Let $\{\mathbf{x}_{c},\mathbf{r}_{c},\mathbf{s},\mathbf{o},\mathbf{c}\}$, $\{\mathbf{x}_{t},\mathbf{r}_{t},\mathbf{s},\mathbf{o},\mathbf{c}\}$, and $\{\mathbf{A}_{t},\mathbf{E}_{t}\}$ denote a Gaussian in $\mathbf{C}_{k}$, the ones in $\mathbf{D}_{t}$, and the corresponding transformation matrix, respectively. It can be written as,
$\mathbf{x}_{t}=\mathbf{A}_{t}\mathbf{x}_{c}+\mathbf{E}_{t},\qquad\mathbf{r}_{t}=\mathbf{A}_{t}\mathbf{r}_{c}.\vspace{-2mm}$ (8)
Kindly refer to Shape-of-Motion [79] for details.
Interpolation with Exposure Time Parameters. To get continuous dynamic Gaussians $\{\mathbf{D}_{t,i}\}_{i=1}^{N}$, one straightforward way is to deploy a series of learnable Gaussian or deformation parameters, but it is unstable to optimize. With the explicit object motion representation in Eq. 8, $\mathbf{D}_{t,i}$ can be calculated by interpolating between the ones at the nearest integer timestamps, i.e.,
$\small\begin{split}\mathbf{D}_{t,i}=\mathbf{w}_{t,i}\odot\mathbf{D}_{t-1}+(1-\mathbf{w}_{t,i})\odot\mathbf{D}_{t},\text{ }{i}\in[1,N/2],\\ \mathbf{D}_{t,i}=(1-\mathbf{w}_{t,i})\odot\mathbf{D}_{t}+\mathbf{w}_{t,i}\odot\mathbf{D}_{t+1},\text{ }{i}\in[N/2,N].\vspace{-1mm}\end{split}$ (9)
$\mathbf{w}_{t,i}$ is the normalized time interval between $\mathbf{D}_{t,i}$ and $\mathbf{D}_{t,N/2}$. Thus, the problem is transformed to estimate $\mathbf{w}_{t,i}$. In the implementation, we can estimate the one at exposure start and end (i.e., $\mathbf{w}_{t,1}$ and $\mathbf{w}_{t,N}$) and then interpolate between them to get the $i$-th intermediate one $\mathbf{w}_{t,i}$, i.e.,
$\mathbf{w}_{t,i}=(1-\frac{i-1}{N-1})\odot\mathbf{w}_{t,1}+\frac{i-1}{N-1}\odot\mathbf{w}_{t,N}.\vspace{-2mm}$ (10)
As the object motion within the exposure can be regarded as uniform, the absolute value of $\mathbf{w}_{t,1}$ and $\mathbf{w}_{t,N}$ are equal, which is half the exposure time $\mathbf{w}_{t}$. Thus, Eq. 10 can be re-written as,
$\mathbf{w}_{t,i}=(1-\frac{i-1}{N-1})\odot\frac{\mathbf{w}_{t}}{2}+\frac{i-1}{N-1}\odot(-\frac{\mathbf{w}_{t}}{2}).\vspace{-1mm}$ (11)
Thus, we set learnable parameters $\mathbf{w}_{t}$ in Deblur4DGS for continuous dynamic Gaussians estimation within the exposure time. The canonical Gaussians, Gaussian deformation modules, and $\mathbf{w}_{t}$ are jointly optimized. The reconstruction loss for dynamic areas is similar to Eq. 6.
Figure 3: Illustration of artifacts in dynamic areas. After optimization with Eq. 4, static areas of $\hat{\mathbf{I}}_{t,i}$ are sharp while dynamic areas exist notable artifacts, as shown in the yellow and red boxes in (b), respectively. The proposed regularization terms $\mathcal{L}_{reg}$ (see Eq. 15) improve the visual effect, as shown in (c).
4.3 Regularization Terms
After optimization with Eq. 4, static areas of $\hat{\mathbf{I}}_{t,i}$ are sharp while dynamic areas can with notable artifacts, as shown in the yellow and red boxes of Fig. 3(b) respectively. The reasons are below. (1) Note that multiple solutions exist for the model to fulfill Eq. 4. The most ideal one is that every $\hat{\mathbf{I}}_{t,i}$ is sharp, and the most trivial one is that every $\hat{\mathbf{I}}_{t,i}$ is as blurry as $\mathbf{B}_{t}$. (2) As static areas are consistent across the entire video, the model tends to learn the underlying sharp representation for inter-frame consistency. In other words, the inter-frame consistency implicitly regularizes model optimization. To further validate this, we conduct an experiment that removes the inter-frame consistency by reducing the number of frames to one. In such a case, the static areas are blurry after optimization with Eq. 4, which supports our confirmation. (3) Compared to static areas, the inter-frame consistency in dynamic ones is weaker due to object motion. It may provide insufficient regularization to guide sharp representation learning, thus leading to results with artifacts. To avoid this, we introduce some regularization terms, including exposure regularization $\mathcal{L}_{e}$, multi-frame consistency regularization $\mathcal{L}_{mfc}$, and multi-resolution consistency regularization $\mathcal{L}_{mrc}$.
First, the continuous dynamic Gaussians $\{\mathbf{D}_{t,i}\}_{i=1}^{N}$ should not be the same. In other words, the value of exposure time parameters $\mathbf{w}_{t}$ should not be too small. If $\mathbf{w}_{t}$ is too small, $\mathbf{D}_{t,i}$ is nearly the same as $\mathbf{D}_{t}$, leading to trivial solutions. We constrain $\mathbf{w}_{t}$ by $\mathcal{L}_{e}$, as,
$\mathcal{L}_{e}=\texttt{max}(0,\mathbf{\epsilon}-\mathbf{w}_{t}).\vspace{-1mm}$ (12)
max is the maximum function and $\mathbf{\epsilon}$ is a threshold.
Second, despite different motions, the content of multiple frames within exposure time should be similar. We utilize $\mathcal{L}_{fc}$ to constrain consistency between neighbor frames, and that between each frame and the first frame, i.e.,
$\begin{split}\mathcal{L}_{mfc}=\frac{1}{N-1}&\sum_{i=2}^{N}(\left\|\mathbf{M}_{t,i}\odot(\hat{\mathbf{I}}_{t,{i-1\rightarrow i}}-\hat{\mathbf{I}}_{t,i})\right\|_{1}\\ &\;\;\>+\left\|\mathbf{M}_{t,1}\odot(\hat{\mathbf{I}}_{t,{i\rightarrow 1}}-\hat{\mathbf{I}}_{t,1})\right\|_{1}).\vspace{-2mm}\end{split}$ (13)
$\hat{\mathbf{I}}_{t,{i-1\rightarrow i}}$ and $\hat{\mathbf{I}}_{t,{i\rightarrow 1}}$ are obtained by aligning $\hat{\mathbf{I}}_{t,{i-1}}$ to $\hat{\mathbf{I}}_{t,i}$ and aligning $\hat{\mathbf{I}}_{t,i}$ to $\hat{\mathbf{I}}_{t,1}$ with a pre-trained optical flow network [71], respectively. $\mathbf{M}_{t,i}$ and $\mathbf{M}_{t,1}$ are masks that indicate dynamic areas in $\hat{\mathbf{I}}_{t,i}$ and $\hat{\mathbf{I}}_{t,1}$, respectively.
Third, the blur in the lower resolution is lower level and is easier to remove [28, 75], thus the artifacts are less in models trained with down-sampled blurry video. Taking advantage of this, we impose $\mathcal{L}_{mrc}$ to assist the optimization of high-resolution models with results from low-resolution models. It can be written as,
$\mathcal{L}_{mrc}=||(\mathbf{M}_{t,i})_{\downarrow}\odot((\hat{\mathbf{I}}_{t,i})_{\downarrow}-\texttt{sg}(\hat{\mathbf{I}}_{t,i}^{l}))||_{1}.\vspace{-2mm}$ (14)
$\hat{\mathbf{I}}_{t,i}^{l}$ is the rendered sharp image from the low-resolution model, which is pre-trained by taking the down-sampled video as supervision. $(\cdot)_{\downarrow}$ is an image down-sampling operation. sg is the stop-gradient operation.
[…]
Figure 4: Visual comparisons of novel-view synthesis on the $720\times 1080$ images. Our method produces more photo-realistic details in both static and dynamic areas, as marked with yellow and red boxes respectively.
Overall, the proposed regularization terms $\mathcal{L}_{reg}$ can be denoted as,
$\mathcal{L}_{reg}=\lambda_{e}\mathcal{L}_{e}+\lambda_{mfc}\mathcal{L}_{mfc}+\lambda_{mrc}\mathcal{L}_{mrc}.\vspace{-1mm}$ (15)
$\lambda_{e}$, $\lambda_{mfc}$, and $\lambda_{mrc}$ are set to 0.1, 2, 1, respectively. Besides, following Shape-of-Motion [79], we also use some other regularization terms $\mathcal{L}_{oth}$ to help reconstruct 3D motion better, and the details can be seen in the Suppl.
4.4 Application to Multiple Tasks
The blurry videos suffer from not only motion blur, but also low frame rates and scene shake generally. Beyond novel-view synthesis, Deblur4DGS can adjust the camera poses and timestamps to address these problems, achieving video deblurring, frame interpolation, and video stabilization. First, when inputting camera poses of the blurry video, Deblur4DGS can render corresponding deblurring results. Second, when feeding the interpolated camera poses and timestamps, Deblur4DGS can produce frame-interpolated results. Third, Deblur4DGS can render a more stable video with the smoothed camera poses as inputs.
5 Experiments
5.1 Experimental Settings
Datasets. We evaluate methods on 6 scenes with significant motion blur from Stereo Blur Dataset [107], where each one contains blurry stereo videos and the corresponding sharp ones. For each scene, we extract 24 frames from the original video and use COLMAP [64] to calibrate the camera extrinsic and intrinsic parameters, following DybluRF [72]. Note that DybluRF [72] conducts experiments on ×$\times 2.5$ down-sampled data. We employ both ×$\times 2.5$ down-sampled ones (i.e., $288\times 512$) and the original ones (i.e., $720\times 1080$).
Training Configurations. We use the blurry image sequences from the left camera for training. For stable optimization, we pre-train the camera motion predictor and static Gaussians $\mathbf{S}$ for $200$ epochs. After that, we jointly optimize the camera motion predictor, $\mathbf{S}$, canonical dynamic Gaussians $\{\mathbf{C}_{k}\}_{k=1}^{K}$, deformable operation $\mathcal{F}$ and exposure time parameters $\{\mathbf{w}_{t}\}_{t=1}^{T}$ for $200$ epochs. The learning rate for the camera motion predictor is set to $5\times 10^{-4}$ and decayed to $1\times 10^{-5}$. The learning rate for $\{\mathbf{w}_{t}\}_{t=1}^{T}$ is set to $1\times 10^{-1}$ and decayed to $1\times 10^{-5}$. The learning rate for $\mathbf{S}$, $\{\mathbf{C}_{k}\}_{k=1}^{K}$ and $\mathcal{F}$ are the same as Shape-of-Motion [79]. $N$ is set to 11. $L$ and $H$ are set to 5 and 3 respectively. $\mathbf{\epsilon}$ is set to $1.0$. All experiments are conducted with PyTorch [57] on one Nvidia GeForce RTX A6000 GPU.
Table 1: Quantitative comparisons for novel view synthesis. $\uparrow$ denotes the higher metric the better, and $\downarrow$ denotes the lower one the better. The best is boldfaced, and the second is underlined.
Methods
[…]
$288\times 512$
[…]
$720\times 1080$
DeformableGS [90] 17.82 / 0.722 / 0.292 17.12 / 0.751 / 0.363
4DGaussians [84] 21.94 / 0.833 / 0.180 21.69 / 0.852 / 0.238
E-D3DGS [1] 22.85 / 0.843 / 0.175 22.38 / 0.854 / 0.240
DyBluRF [72] 27.09 / 0.919 / 0.090 26.46 / 0.913 / 0.142
Shape-of-Motion [79] 27.89 / 0.928 / 0.116 27.49 / 0.922 / 0.192
Deblur4DGS (Ours) 28.92 / 0.949 / 0.060 28.88 / 0.947 / 0.098
Table 2: Quantitative results that pre-process the blurry video with an image or video deblurring method before 4D reconstruction. ‘None’ denotes no image or video deblurring method being used.
Reconstruction
Methods
Deblurring
Methods
[…]
$288\times 512$
[…]
$720\times 1080$
E-D3DGS [1] None 22.85 / 0.843 / 0.175 22.38 / 0.854 / 0.240
Restormer [92] 23.55 / 0.854 / 0.142 22.77 / 0.867 / 0.186
DSTNet [56] 23.21 / 0.863 / 0.132 22.80 / 0.865 / 0.177
BSSTNet [95] 23.28 / 0.857 / 0.128 23.08 / 0.876 / 0.163
Shape-of-Motion [79] None 27.89 / 0.928 / 0.116 27.49 / 0.922 / 0.192
Restormer [92] 28.39 / 0.941 / 0.080 28.25 / 0.940 / 0.117
DSTNet [56] 28.07 / 0.938 / 0.072 28.22 / 0.941 / 0.104
BSSTNet [95] 28.34 / 0.940 / 0.079 28.24 / 0.943 / 0.100
Deblur4DGS (Ours) None 28.92 / 0.949 / 0.060 28.88 / 0.947 / 0.098
Evaluation Configurations. We evaluate methods on novel-view synthesis, deblurring, frame interpolation, and video stabilization tasks. To evaluate novel-view synthesis and deblurring performance, we use the sharp image sequences of the right and left cameras as ground truth respectively. Due to calibrated camera parameters are not absolutely accurate, the rendering results may be slightly spatially misaligned with the ground truth. We adopt the aligned PSNR [22], SSIM [81] and LPIPS [97] as evaluation metrics, following similar works [3, 12, 85]. Specifically, we align rendering results to the ground truth by a pre-trained optical flow network (i.e., PWC-Net [71]) and then calculate metrics between the aligned results and the ground truth. Simultaneously, we report results that calculate metrics between the original rendering images and the ground truth in the Suppl. As there is no ground truth for frame interpolation and video stabilization tasks, we evaluate method performance with recent no-reference metrics, i.e., CLIPIQA [77] and MUSIQ [26].
Table 3: Quantitative comparisons for deblurring. The first category is supervised image or video deblurring methods, and the second is 4D reconstruction-based ones. The best one in each category is boldfaced, and the second one is underlined.
Methods
[…]
$288\times 512$
[…]
$720\times 1080$
Restormer [92] 33.33 / 0.974 / 0.043 32.43 / 0.962 / 0.064
DSTNet [56] 33.13 / 0.971 / 0.039 32.42 / 0.962 / 0.059
BSSTNet [95] 34.13 / 0.980 / 0.027 33.63 / 0.973 / 0.039
DeformableGS [90] 30.92 / 0.949 / 0.103 29.11 / 0.926 / 0.191
4DGaussians [84] 31.25 / 0.949 / 0.097 29.95 / 0.932 / 0.169
E-D3DGS [1] 31.28 / 0.949 / 0.098 29.98 / 0.931 / 0.170
DyBluRF [72] 29.87 / 0.947 / 0.054 29.23 / 0.943 / 0.105
Shape-of-Motion [79] 29.83 / 0.946 / 0.107 28.93 / 0.932 / 0.190
Deblur4DGS (Ours) 32.55 / 0.975 / 0.041 31.60 / 0.965 / 0.084
5.2 Comparison with State-of-the-Art Methods
We compare Deblur4DGS with 5 state-of-the-art 4D reconstruction methods (i.e., DeformableGS [90], 4DGaussians [84], E-D3DGS [1], DyBluRF [72], and Shape-of-Motion [79]) on two resolution (i.e., $228\times 512$ and $720\times 1080$) datasets. We input poses of right and left cameras to produce novel-view synthesis and deblurring results, respectively. We linearly interpolate the left camera poses and timestamps to generate ×$\times 16$ frame interpolation results. We employ a Gaussian filter to smooth left camera poses for video stabilization, following  [59]. The results for novel-view synthesis, deblurring, frame interpolation, and video stabilization are summarized in Tab. A, Tab. B, Tab. 4, and Tab. 5, respectively. Besides, rendering speed comparisons on $720\times 1080$ images are reported in Fig. 1.
Novel-view synthesis Comparisons. As shown in Tab. A, among all competing methods, DyBluRF [72] gets the best LPIPS score, while Shape-of-Motion [79] achieves the best PSNR and SSIM scores. Benefiting the explicit 3D representation manner and elaborate design for processing motion blur, our Deblur4DGS obtains all the best scores on both datasets. In addition, Deblur4DGS removes blur more clearly and produces more photo-realistic details in both static and dynamic areas, as shown in Fig. 4. More visual comparisons can be seen in the Suppl.
In addition, to further demonstrate the effectiveness of Deblur4DGS, we first pre-process the blurry videos with state-of-the-art image [92] or video [56, 95] deblurring methods and then perform 4D reconstruction. The results are summarized in Tab. 2. Compared with reconstruction from blurry videos, the incorporation of deblurring models improves performance. This is because the deblurring models remove some blur, facilitating sharp scene reconstruction. However, as the image and video deblurring methods cannot perceive 3D structure and maintain scene geometric consistency, the scene reconstruction results are still unsatisfactory. In contrast, Deblur4DGS jointly reconstructs scene geometry and processes motion blur in 3D space, achieving better scene reconstruction results.
Table 4: Quantitative comparisons for frame interpolation. The first category is supervised video frame interpolation methods, and the second is 4D reconstruction-based ones.
Methods
[…]
$288\times 512$
[…]
$720\times 1080$
RIFE [21] 0.179 / 44.090 0.147 / 32.637
EMAVFI [94] 0.181 / 42.773 0.170 / 32.564
VIDUE [66] 0.267 / 60.854 0.185 / 49.560
DeformableGS [90] 0.172 / 41.603 0.167 / 31.205
4DGaussians [84] 0.177 / 44.132 0.177 / 33.836
E-D3DGS [1] 0.178 / 43.250 0.181 / 32.945
DyBluRF [72] 0.166 / 51.513 0.132 / 36.768
Shape-of-Motion [79] 0.176 / 43.020 0.202 / 33.817
Deblur4DGS (Ours) 0.200 / 54.043 0.213 / 42.139
Table 5: Quantitative comparisons for video stabilization. The first category is 2D video stabilization methods, and the second is 4D reconstruction-based ones.
Methods
[…]
$288\times 512$
[…]
$720\times 1080$
MeshFlow [47] 0.140 / 38.289 0.116 / 34.140
NNDVS [100] 0.186 / 45.613 0.137 / 33.269
DeformableGS [90] 0.170 / 40.846 0.168 / 30.971
4DGaussians [84] 0.145 / 44.673 0.172 / 34.416
E-D3DGS [1] 0.181 / 44.364 0.182 / 33.828
DyBluRF [72] 0.164 / 53.115 0.132 / 37.556
Shape-of-Motion [79] 0.174 / 43.293 0.202 / 33.750
Deblur4DGS (Ours) 0.199 / 54.299 0.211 / 42.241
Deblurring Comparisons. Apart from 4D reconstruction-based methods, we additionally compare with some state-of-the-art image [92] and video [56, 95] deblurring ones. As shown in Tab. B, Deblur4DGS obtains better results than 4D reconstruction-based methods and comparable ones to the deblurring-specific (i.e., Restormer [92], DSTNet [56], and BSSTNet [95]) methods. Compared with the former, Deblur4DGS better reconstructs the scene, thus performing better. Note that the latter ones are trained on large paired data in a fully supervised manner while Deblur4DGS is optimized only with the given blurry video in a self-supervised manner. Although the large data prior makes these methods perform better, Deblur4DGS is more convenient to use without paired data collection.
Frame Interpolation Comparisons. We compare Deblur4DGS with 4D reconstruction-based methods and some state-of-the-art supervised video frame interpolation ones (i.e., RIFE [21], EMAVFI [94], and VIDUE [66]), as shown in Tab. 4. Deform4DGS generally performs satisfactorily. Note that VIDUE [66] is specially designed and trained with large paired data in a supervised manner for joint deblurring and fame Interpolation, thus achieving better results.
Video Stabilization Comparisons. The results are summarized in Tab. 5. Benefiting from the better geometry reconstruction, Deblur4DGS achieves pleasant scores compared with both 2D video stabilization methods (i.e., MeshFlow [47] and NNDVS [100]) and 4D reconstruction-based ones, which further demonstrates its effectiveness.
Figure 5: Effect about the estimation of continuous camera poses (ECP) and dynamic Gaussians (EDG). Static areas are marked with yellow boxes while dynamic ones are marked with red boxes. Without ECP and EDG, the results are almost as blurry as the input frame, as shown in (b). Only with ECP, static areas are sharp but may lead to visual artifacts in dynamic areas, as shown in (c). Further introducing EDG improves this, as shown in (d).
6 Ablation Study
We first discuss the estimation of continuous camera poses (ECP) and dynamic Gaussians (EDG). Then, we validate the effect of regularization terms, as well as the blur-aware variable (BAV) canonical Gaussians.
6.1 Effect of ECP and EDG
ECP and EDG are introduced to process camera motion blur and object motion blur, respectively. Visual comparisons are shown in Fig. 5, and quantitative results are in the Suppl. First, without ECP and EDG, the results are almost as blurry as the input frame, as shown in Fig. 5(b). Second, only with ECP, the static areas are sharp but may lead to visual artifacts in dynamic areas, as shown in Fig. 5(c). It is because ECP cannot simulate the object movement. Third, we further introduce EDG to simulate that produces visually pleasant results in both areas, as shown in Fig. 5(d).
6.2 Effect of Regularization Terms
The effect of exposure regularization $\mathcal{L}_{e}$ (see Eq. 12, multi-frame consistency regularization $\mathcal{L}_{mfc}$ (see Eq. 13) and multi-resolution consistency regularization $\mathcal{L}_{mrc}$ (see Eq. 14) are shown in Tab. D. Without these regularization terms, the performance drops significantly and the visual results in Fig. 3 show that dynamic areas are with notable artifacts. By regularizing the object motion within the exposure time distinguished, $\mathcal{L}_{e}$ alleviates the trivial solutions and improves performance. Besides, $\mathcal{L}_{mfc}$ and $\mathcal{L}_{mrc}$ additionally regularize multi-frame and multi-resolution consistency respectively, helping to alleviate artifacts. Their combinations perform best.
6.3 Effect of BAV Canonical Gaussians.
The effect of BAV canonical Gaussians is summarized in Tab. E. First, selecting a single canonical Gaussians across the entire video (i.e., None) leads to poor performance, due to the challenge of modeling large object motion. Second, selecting variable canonical Gaussians uniformly (i.e., w/o Blur-Aware) alleviates the challenge, thus leading to performance improvement. We also experiment with an optical flow-based strategy [67] to select the canonical Gaussians, and it achieves a similar performance to the uniform selection. This may be due to the inaccurate estimation of optical flow from blurry images. Third, our blur-aware selection is better, as the canonical Gaussians corresponding to the sharper frame help blur removal.
Table 6: Ablation studies on regularization terms (see Eq. 15).
[…]
$\times$ $\times$ $\times$ 28.69 / 0.947 / 0.063
$\times$ $\checkmark$ $\checkmark$ 28.86 / 0.949 / 0.060
$\checkmark$ $\times$ $\times$ 28.73 / 0.947 / 0.060
$\checkmark$ $\checkmark$ $\times$ 28.87 / 0.948 / 0.060
$\checkmark$ $\times$ $\checkmark$ 28.86 / 0.948 / 0.060
$\checkmark$ $\checkmark$ $\checkmark$ 28.92 / 0.949 / 0.060
Table 7: Effect of blur-aware variable (BAV) canonical Gaussians. ‘None’ denotes selecting a single one across the entire video.
[…]
None 28.67 / 0.946 / 0.065
w/o Blur-Aware 28.85 / 0.949 / 0.062
Ours 28.92 / 0.949 / 0.060
7 Conclusions
In this work, we propose Deblur4DGS, the first 4D Gaussian Splatting framework to reconstruct a high-quality 4D model from blurry monocular video. In particular, with the explicit motion trajectory modeling based on 3D Gaussian Splatting, we propose to transform the challenging continuous dynamic representation estimation within an exposure time into the exposure time estimation, where a series of regularizations are suggested to tackle the under-constrained optimization. Besides, a blur-aware variable canonical Gaussians is present to represent objects with large motion better. Beyond novel-view synthesis, Deblur4DGS can improve blurry video quality from multiple perspectives, including deblurring, frame interpolation, and video stabilization. Extensive results show Deblur4DGS outperforms state-of-the-art 4D reconstruction methods.
[…]
Supplementary Material
The content of the supplementary material involves:
  • Structure of camera motion predictor in Sec. A.
  • Other regularization terms in Sec. B.
  • More result comparisons in Sec. C.
  • Effect of ECP and EDG in Sec. D.
  • Effect of regularization terms in Sec. E.
  • Effect of BAV canonical Gaussians in Sec. F.
A Structure of Camera Motion Predictor
The structure of camera motion predictor is provided in Fig. A. It first embeds the camera pose $\mathbf{P}_{t}$ to a higher dimensional space, following [52]. Then, three FC blocks are stacked, each consisting of an FC layer followed by a ReLU operation. Finally, we deploy two heads to predict the camera pose at exposure start and end (i.e., $\mathbf{P}_{t,1}$ and $\mathbf{P}_{t,N}$), respectively.
B Other Regularization Terms
Following Shape-of-Motion [79], we use some other regularization terms $\mathcal{L}_{oth}$ to help reconstruct 3D motion better, including mask regularization $\mathcal{L}_{mask}$, 2D tracks regularization $\mathcal{L}_{track}$ and distance-preserving regularization $\mathcal{L}_{rigid}$.
Specifically, we render the masks within the exposure time $\{\hat{\mathbf{M}}_{t,i}\}_{i=1}^{N}$ to indicate dynamic areas. To supervise the training, we synthesize the mask $\hat{\mathbf{M}}_{t}^{B}$ for the synthetic blurry image $\hat{\mathbf{B}}_{t}$ as,
$\hat{\mathbf{M}}_{t}^{B}(u,v)=\texttt{max}\{\hat{\mathbf{M}}_{t,1}(u,v),\hat{\mathbf{M}}_{t,2}(u,v),...,\hat{\mathbf{M}}_{t,N}(u,v)\}.$ (A)
$(u,v)$ is the pixel location. The mask regularization $\mathcal{L}_{mask}$ can be written as,
$\mathcal{L}_{mask}=\mathcal{L}_{1}(\hat{\mathbf{M}}_{t}^{B},\mathbf{M}_{t}),$ (B)
where $\mathbf{M}_{t}$ is the mask that obtained by applying SAM2 [61] to the ground truth blurry frame. Besides, we render the 2D tracks $\hat{\mathbf{U}}_{t\rightarrow t^{\prime}}$ from a pair of randomly sampled query time $t$ and target time $t^{\prime}$. We supervise it by the lifted long-range 2D tracks $\mathbf{U}_{t\rightarrow t^{\prime}}$ that are extracted from TAPIR [9], i.e.,
$\mathcal{L}_{track}=\mathcal{L}_{1}(\hat{\mathbf{U}}_{t\rightarrow t^{\prime}},\mathbf{U}_{t\rightarrow t^{\prime}}).$ (C)
Finally, we enforce a distance-preserving loss $\mathcal{L}_{rigid}$ between randomly sampled dynamic Gaussians and their $J$-nearest neighbors. Let $\mathbf{x}_{t}$ and $\mathbf{x}_{t^{\prime}}$ denote the position of a Gaussian at time $t$ and $t^{\prime}$, and $\mathcal{C}_{J}(\mathbf{x}_{t})$ denote the set of $J$-nearest neighbors of $\mathbf{x}_{t}$. $\mathcal{L}_{rigid}$ can be written as,
$\mathcal{L}_{rigid}=||\texttt{dist}(\hat{\mathbf{x}}_{t},\mathcal{C}_{J}(\mathbf{x}_{t}))-\texttt{dist}(\hat{\mathbf{x}}_{t^{\prime}},\mathcal{C}_{J}(\mathbf{x}_{t^{\prime}}))||_{2}^{2}.$ (D)
dist measures the Euclidean distance.
Figure A: Structure of camera motion predictor.
Overall, $\mathcal{L}_{oth}$ can be written as,
$\mathcal{L}_{oth}=\lambda_{mask}\mathcal{L}_{mask}+\lambda_{track}\mathcal{L}_{track}+\lambda_{rigid}\mathcal{L}_{rigid}.$ (E)
$\lambda_{mask}$, $\lambda_{track}$ and $\lambda_{rigid}$ are set to 1, 2, and 2, respectively.
C More Result Comparisons
Due to calibrated camera parameters are not absolutely accurate, the rendering results may be slightly spatially misaligned with the ground truth. In the main text, we calculate the metrics (i.e., PSNR [22], SSIM [81] and LIPIS [97]) between the aligned rendering results and the ground truth. Here, we provide the novel-view synthesis and deblurring comparisons that directly calculate metrics between the rendering results and the ground truth, as shown in Tab. A and Tab. B respectively. It shows that Deblur4DGS still outperforms state-of-the-art 4D reconstruction-based methods. More visual results for novel view synthesis in Fig. B show that our method produces more photo-realistic details. Moreover, visual results for deblurring in Fig. C show that compared with 4D reconstruction methods, Deblur4DGS produces sharper contents and fewer visual artifacts. Furthermore, the frame interpolation and video stabilization results, as well as some novel view synthesis videos are shown in the site.
D Effect of ECP and EDG
The visual results are provided in the main text. Here we provide the quantitative results in Tab. C. Without ECP and EDG, the reconstruction results are unsatisfactory due to motion blur. ECP and EDG enable processing of the camera motion blur and object motion blur respectively, thus leading to performance improvement.
E Effect of regularization terms
We validate the effect of exposure regularization $\mathcal{L}_{e}$, multi-frame consistency regularization $\mathcal{L}_{mfc}$ and multi-resolution consistency regularization $\mathcal{L}_{mrc}$. In the main text, we provide the results on $288\times 512$ data. Here, we provide that on $720\times 1080$ data in Tab. D. The results are consistent for both data. $\mathcal{L}_{e}$ regularizes the object motion within the exposure time distinguished, alleviating the trivial solutions and improving performance. $\mathcal{L}_{mfc}$ and $\mathcal{L}_{mrc}$ regularize multi-frame and multi-resolution consistency respectively, helping to alleviate artifacts. Their combinations perform best.
F Effect of BAV canonical Gaussians
Tab. E additionally provides the quantitative results on $720\times 1080$ data, which are consistent with the ones on $288\times 512$ data. Compared with selecting a single one across the entire video (i.e., None), selecting variable canonical Gaussians uniformly (i.e., w/o Blur-Aware) alleviates the challenge of modeling large object motion, thus leading to performance improvement. Our method selects the canonical Gaussians corresponding to the sharper frame, which is beneficial to blur removal and performs better.
Table A: Quantitative comparisons for novel view synthesis. The metrics are directly calculated between rendering results and ground truth. The best is boldfaced, and the second is underlined.
Methods
[…]
$288\times 512$
[…]
$720\times 1080$
DeformableGS [90] 16.32 / 0.644 / 0.356 16.06 / 0.711 / 0.424
4DGaussians [84] 20.58 / 0.795 / 0.186 20.31 / 0.821 / 0.258
E-D3DGS [1] 20.99 / 0.804 / 0.179 20.95 / 0.829 / 0.257
DyBluRF [72] 23.39 / 0.823 / 0.088 22.59 / 0.815 / 0.168
Shape-of-Motion [79] 26.47 / 0.908 / 0.104 25.76 / 0.896 / 0.195
Deblur4DGS (Ours) 26.69 / 0.918 / 0.052 25.93 / 0.904 / 0.102
Table B: Quantitative comparisons for deblurring. The metrics are directly calculated between rendering results and ground truth. The first category is supervised image or video deblurring methods, and the second is 4D reconstruction-based ones.
Methods
[…]
$288\times 512$
[…]
$720\times 1080$
Restormer [92] 36.74 / 0.987 / 0.019 35.97 / 0.981 / 0.028
DSTNet [56] 36.49 / 0.986 / 0.015 35.24 / 0.977 / 0.027
BSSTNet [95] 36.88 / 0.988 / 0.013 35.87 / 0.983 / 0.021
DeformableGS [90] 30.97 / 0.950 / 0.096 28.95 / 0.924 / 0.189
4DGaussians [84] 31.30 / 0.951 / 0.090 29.76 / 0.929 / 0.167
E-D3DGS [1] 31.34 / 0.951 / 0.091 29.74 / 0.929 / 0.170
DyBluRF [72] 29.32 / 0.942 / 0.049 28.67 / 0.936 / 0.101
Shape-of-Motion [79] 29.51 / 0.945 / 0.095 28.35 / 0.927 / 0.187
Deblur4DGS (Ours) 31.77 / 0.972 / 0.033 30.48 / 0.956 / 0.081
Table C: Effect about the estimation of continuous camera poses (ECP) and dynamic Gaussians (EDG).
ECP EDG
[…]
$288\times 512$
[…]
$720\times 1080$
$\times$ $\times$ 28.00 / 0.929 / 0.119 27.70 / 0.923 / 0.191
$\checkmark$ $\times$ 28.86 / 0.948 / 0.062 28.68 / 0.945 / 0.100
$\checkmark$ $\checkmark$ 28.92 / 0.949 / 0.060 28.88 / 0.947 / 0.098
Table D: Ablation studies on regularization terms.
$\mathcal{L}_{e}$ $\mathcal{L}_{mfc}$ $\mathcal{L}_{mrc}$
[…]
$288\times 512$
[…]
$720\times 1080$
$\times$ $\times$ $\times$ 28.69 / 0.947 / 0.063 28.70 / 0.946 / 0.099
$\times$ $\checkmark$ $\checkmark$ 28.86 / 0.949 / 0.060 28.86 / 0.947 / 0.098
$\checkmark$ $\times$ $\times$ 28.73 / 0.947 / 0.060 28.74 / 0.945 / 0.099
$\checkmark$ $\checkmark$ $\times$ 28.87 / 0.948 / 0.060 28.78 / 0.946 / 0.098
$\checkmark$ $\times$ $\checkmark$ 28.86 / 0.948 / 0.060 28.79 / 0.946 / 0.099
$\checkmark$ $\checkmark$ $\checkmark$ 28.92 / 0.949 / 0.060 28.88 / 0.947 / 0.098
Table E: Effect of blur-aware variable (BAV) canonical Gaussians. ‘None’ denotes selecting a single one across the entire video.
Methods
[…]
$288\times 512$
[…]
$720\times 1080$
None 28.67 / 0.946 / 0.065 28.66 / 0.945 / 0.104
w/o Blur-Aware 28.85 / 0.949 / 0.062 28.83 / 0.947 / 0.098
Ours 28.92 / 0.949 / 0.060 28.88 / 0.947 / 0.098
[…]
Figure B: Visual comparisons of novel-view synthesis on the $720\times 1080$ images. Our method produces more photo-realistic details in both static and dynamic areas, as marked with yellow and red boxes respectively.
[…]
Figure C: Visual comparisons of deblurring on the $720\times 1080$ images. Compared with 4D reconstruction-based methods, Deblur4DGS produces sharper contents and fewer artifacts in both static and dynamic areas, as marked with yellow and red boxes respectively.