### v3

# Deblur4DGS: 4D Gaussian Splatting from Blurry Monocular Videos

Abstract
Recent 4D reconstruction methods have yielded impressive results but rely on sharp videos as supervision. However, motion blur often occurs in videos due to camera shake and object movement, while existing methods render blurry results when using such videos for reconstructing 4D models. Although a few approaches attempted to address the problem, they struggled to produce high-quality results, due to the inaccuracy in estimating continuous dynamic representations within the exposure time. Encouraged by recent works in 3D motion trajectory modeling using 3D Gaussian Splatting (3DGS), we take 3DGS as the scene representation manner, and propose Deblur4DGS to reconstruct a high-quality 4D model from blurry monocular video. Specifically, we transform continuous dynamic representations estimation within an exposure time into the exposure time estimation. Moreover, we introduce the exposure regularization term, multi-frame, and multi-resolution consistency regularization term to avoid trivial solutions. Furthermore, to better represent objects with large motion, we suggest blur-aware variable canonical Gaussians. Beyond novel-view synthesis, Deblur4DGS can be applied to improve blurry video from multiple perspectives, including deblurring, frame interpolation, and video stabilization. Extensive experiments in both synthetic and real-world data on the above four tasks show that Deblur4DGS outperforms state-of-the-art 4D reconstruction methods. The codes are available at https://github.com/ZcsrenlongZ/Deblur4DGS.
Introduction
Substantial efforts have been made for 4D reconstruction, which has extensive applications in augmented reality and virtual reality. To model static scenes, Neural Radiance Field (NeRF) (Mildenhall et al. 2021) and 3D Gaussian Splatting (3DGS) (Kerbl et al. 2023) propose implicit neural representation manner and explicit Gaussian ellipsoids one, respectively. To model dynamic objects, implicit neural fields (Zhu et al. 2024; Yang et al. 2024; Wu et al. 2024; Yan et al. 2023) and explicit deformation (Duan et al. 2024b; Chu et al. 2024; Katsumata et al. 2024; Lin et al. 2024; Li et al. 2024; Wang et al. 2024) are suggested for motion representation. While achieving great progress, most methods rely on synchronized multi-view videos. They yield unsatisfactory results when applied to monocular video, where dynamic objects are only observed once at each timestamp. To alleviate the under-constrained nature of the problem, recent studies have introduced data-driven priors, such as depth maps (Lee et al. 2023c; Yang et al. ), optical flows (Gao et al. 2024; Zhu et al. 2024), tracks (Seidenschwarz and et al. 2024; Lei et al. 2024), and generative models (Wu et al. 2025; Chu et al. 2024; Zeng et al. 2025) for better 4D reconstruction.
Unfortunately, motion blur often arises due to camera shake and object movement. When reconstructing the 4D scene from the blurry video, the above methods usually render blurry results. The first step to solving this problem is to deal with camera motion blur, which is relatively simple. Some NeRF-based (Lee et al. 2023a; Wang et al. 2023b; Lee et al. 2023b) and 3DGS-based (Zhao et al. 2024; Chen and Liu 2024; Oh et al. 2024) methods have suggested jointly optimizing 3D representation and camera poses within the exposure time by calculating the reconstruction loss between the synthetic blurry images and the input blurry frames. In contrast, the object motion blur is more challenging to address, as the solution has to estimate continuous and sharp dynamic representations within the exposure time to simulate blurry frames.
In this work, we take 3DGS (Kerbl et al. 2023) as the scene representation manner to explore the problem, driven by two main motivations. First, its successful application in 4D reconstruction make this method highly promising. Second, the explicit 3D motion modeling presents an opportunity to simplify the complex continuous dynamic representations estimation within the exposure time into exposure time estimation, avoiding complex extra motion modeling in DyBluRF (Sun et al. 2024a; Bui and et al. 2023). Once the exposure time is estimated, continuous dynamic representations can be obtained by directly interpolating between representations at the nearest integer timestamps. We note that the concurrent work BARD-GS (Lu et al. 2025) adopts a similar strategy, but they perform unsatisfactorily due to the under-constrained optimization, especially for large object motion.
Specifically, we propose Deblur4DGS, a Gaussian Splatting framework for 4D reconstruction from blurry monocular video. For the static scene, we jointly optimize the camera poses at exposure start and end with static Gaussians. For the dynamic objects, we optimize learnable exposure time parameters and dynamic Gaussians of the integer timestamps, simultaneously. Then continuous camera poses and dynamic Gaussians within exposure time can be obtained by interpolation, and they are used to render continuous sharp frames to calculate the reconstruction loss. Moreover, to avoid trivial solutions, we introduce the exposure regularization term, as well as the multi-frame and multi-resolution consistency regularization terms. Furthermore, existing 4D reconstruction methods generally select Gaussians at a single timestamp as canonical Gaussians. However, it may produce results with missing details in scenes with large motion, especially when processing blurry videos with a low frame rate. To alleviate this issue, we suggest variable canonical Gaussians as time progresses based on the image blur level. Gaussians corresponding to the sharper frame are selected as the canonical ones for better blur removal, and each canonical Gaussian is only used for nearby timestamps to reduce difficulty of modeling large motion.
Blurry videos suffer from not only motion blur, but also low frame rates and scene shake generally. Beyond novel-view synthesis, the optimized Deblur4DGS can be applied to address these problems, achieving deblurring, frame interpolation, and video stabilization. We evaluate Deblur4DGS from all four perspectives. Extensive experiments on both synthetic and real-world data demonstrate that Deblur4DGS outperforms state-of-the-art 4D reconstruction methods quantitatively and qualitatively while maintaining real-time rendering speed. Furthermore, Deblur4DGS has competitive capabilities in comparison with task-specific video processing models trained in a supervised manner.
The main contributions can be summarized as follows:
  • We propose Deblur4DGS, a 4D Gaussian Splatting framework specially designed to reconstruct a high-quality 4D model from blurry monocular video.
  • We propose transforming dynamic representation estimation into exposure time estimation, where a series of regularizations are suggested to tackle under-constrained optimization and blur-aware variable canonical Gaussians is present to better represent dynamic objects.
  • Extensive experiments in synthetic and real-world data show that Deblur4DGS outperforms state-of-the-art 4D reconstruction methods on novel-view synthesis, deblurring, frame interpolation, and video stabilization tasks.
Related Work
Image and Video Deblurring
Deep learning-based image (Ren et al. 2023; Li and et al. 2023; Wang et al. 2022; Zhang et al. 2022, 2024) and video (Zhong et al. 2023b; Pan et al. 2023; Zhong et al. 2023a, 2020; Chan et al. 2022) deblurring methods have been widely explored. Compared to image deblurring methods, video ones leverage temporal clues between consecutive frames for more effective restoration. DSTNet (Pan et al. 2023) develops a deep discriminative spatial and temporal network. BasicVSR++ (Chan et al. 2022) improves feature fusion with second-order feature propagation and flow-guided alignment. BSSTNet (Zhang and et al. 2024) introduces a blur map to sufficiently utilize the entire video, achieving recent state-of-the-art. When reconstructing from a blurry video, pre-processing it with the 2D deblurring method is a straightforward manner. However, 2D deblurring methods cannot perceive 3D structures and maintain scene geometric consistency, leading to unsatisfactory scene reconstruction.
Figure 1: (a) Training of Deblur4DGS. When processing $t$-th frame, we first discretize its exposure time into $N$ timestamps. Then, we estimate continuous camera poses $\{\mathbf{P}_{t,i}\}_{i=1}^{N}$ and dynamic Gaussians $\{\mathbf{D}_{t,i}\}_{i=1}^{N}$ within exposure time. Next, we render each latent sharp image $\hat{\mathbf{I}}_{t,i}$ with the camera pose $\mathbf{P}_{t,i}$, dynamic Gaussians $\mathbf{D}_{t,i}$ and static Gaussians $\mathbf{S}$. Finally, $\{\hat{\mathbf{I}}_{t,i}\}_{i=1}^{N}$ are averaged to obtain the synthetic blurry image $\hat{\mathbf{B}}_{t}$, which is used to calculate the reconstruction loss $\mathcal{L}_{rec}$ with the given blurry frame $\mathbf{B}_{t}$. To regularize the under-constrained optimization, we introduce exposure regularization $\mathcal{L}_{e}$, multi-frame consistency regularization $\mathcal{L}_{mfc}$ and multi-resolution consistency regularization $\mathcal{L}_{mrc}$. (b) Rendering of Deblur4DGS. Deblur4DGS produces the sharp image with user-provided timestamp $t$ and camera pose $\mathbf{P}_{t}$.
3D and 4D Reconstruction
To reconstruct 3D models, NeRF (Mildenhall et al. 2021) and 3DGS (Kerbl et al. 2023) introduce implicit neural representation manner and explicit Gaussian ellipsoids one respectively, where the latter generally achieves better results. To reconstruct 4D models, most works (Somraj et al. 2024; Duan et al. 2024a; Lu et al. 2024; Lin et al. 2024; Li et al. 2024; Sun et al. 2024b; Wu et al. 2024; Yang et al. 2024; Mihajlovic et al. 2024; Wang and et al. 2025) incorporate implicit neural fields and explicit deformation for motion representation. Moreover, to better reconstruct from monocular video, some studies enhance 4D reconstruction with data-driven priors, such as depth maps (Lee et al. 2023c; Yang et al. ), optical flows (Gao et al. 2024; Wang and et al. 2025), tracks (Wang et al. 2024; Seidenschwarz and et al. 2024), and generative models (Wu et al. 2025; Chu et al. 2024). For example, GFlow (Wang and et al. 2025) utilizes only 2D priors to lift a video to a 4D scene. GaussianMarbles (Stearns et al. 2024) reduces the degrees of freedom of each Gaussian.
Note that these methods heavily rely on high-quality sharp videos for supervision and perform poorly when facing blurry inputs. To process camera motion in static areas, recent works (Lee et al. 2023a; Ma et al. 2022; Wang et al. 2023b; Lee et al. 2023b, 2024a; Zhao et al. 2024; Peng et al. 2024a; Lee et al. 2024b; Chen and Liu 2024; Oh et al. 2024) suggest jointly optimizing the scene representation and recovering the camera poses within the exposure time. To process object motion blur in dynamic scenes, DyBluRF (Sun et al. 2024a; Bui and et al. 2023) incorporates object motion blur formation into dynamic model optimization but faces challenges in producing high-quality images and achieving real-time rendering. In this work, with 3DGS as the scene representation manner, we develop Deblur4DGS to reconstruct a high-quality 4D model from a blurry video.
Preliminary
4D Gaussian Splatting
A 3D Gaussian (Kerbl et al. 2023) is parameterized by $\{\mathbf{x},\mathbf{r},\mathbf{s},\mathbf{o},\mathbf{c}\}$, where $\mathbf{x}$ characterizes the center position in the world space, rotation matrix $\mathbf{r}$ and scale matrix $\mathbf{s}$ define the shape, $\alpha$ is opacity, and spherical harmonics (SH) coefficients $\mathbf{c}$ represent the view-dependent color.
4D Gaussian Splatting (4DGS) usually process static and dynamic regions separately. Static regions can be represented by a set of 3D Gaussians, named $\mathbf{S}$. For the dynamic areas, 4DGS generally selects a timestamp (e.g., the first timestamp) and represents the objects by canonical dynamic Gaussians, i.e., $\mathbf{C}$. Then, $\mathbf{C}$ is deformed to other timestamps for motion representation. Denote by $\mathbf{D}_{t}$ the dynamic Gaussians at $t$-th timestamp, it can be written as,
$\mathbf{D}_{t}=\mathcal{F}(\mathbf{C},t;\Theta_{\mathcal{F}}).$ (1)
$\mathcal{F}$ is the deformation operation with parameters $\Theta_{\mathcal{F}}$. The Gaussians for $t$-th timestamp is the union of $\mathbf{S}$ and $\mathbf{D}_{t}$.
Collectively, 4DGS models a scene with static Gaussians $\mathbf{S}$, canonical dynamic Gaussians $\mathbf{C}$, and a deformation operation $\mathcal{F}$. With the provided camera pose, the Gaussians at $t$-th timestamp $\mathbf{D}_{t}$ can be projected into 2D spaces and rasterized to obtain the corresponding image.
Motion Blur Formation
Motion blur occurs due to camera shake and object movement, which can be regarded as the integration of latent sharp images (Nah and et al. 2017), i.e.,
$\mathbf{B}(u,v)=\phi\int_{0}^{\tau}\mathbf{I}_{t}(u,v)dt.$ (2)
$\mathbf{B}\in\mathbb{R}^{H\times W\times 3}$ is the blurry image and $\mathbf{I}_{t}$ is the latent sharp one at $t$-th timestamp. $(u,v)$ is pixel location, $\tau$ is the camera exposure time, and $\phi$ is a normalization factor. To approximate the integral operation, recent works (Zhao et al. 2024; Sun et al. 2024a; Wang et al. 2023b) divide the exposure time into $N$ timestamps and regard the blurry image as the average of $N$ sharp images, i.e.,
$\mathbf{B}(u,v)\approx\frac{1}{N}\sum_{i=0}^{N-1}\mathbf{I}_{i}(u,v).$ (3)
In this work, we reconstruct 4D model from a blurry video by integrating the blur formation into model optimization.
Proposed Method
Let $\{\mathbf{B}_{t}\}_{t=1}^{T}$ and $\{\mathbf{M}_{t}\}_{t=1}^{T}$ denote a blurry video with $T$ timestamps and the corresponding masks indicating dynamic areas (extracted by SAM2 (Ravi et al. 2024)), respectively. As shown in fig. 1(a), when processing $t$-th frame, we first evenly divide its camera exposure time into $N$ timestamps. Then, we estimate continuous camera poses $\{\mathbf{P}_{t,i}\}_{i=1}^{N}$ and dynamic Gaussians $\{\mathbf{D}_{t,i}\}_{i=1}^{N}$ to simulate camera shake and object movement. Next, we render each sharp image $\hat{\mathbf{I}}_{t,i}$ with the corresponding camera pose $\mathbf{P}_{t,i}$, dynamic Gaussians $\mathbf{D}_{t,i}$ and static Gaussians $\mathbf{S}$. After that, we average $\{\hat{\mathbf{I}}_{t,i}\}_{i=1}^{N}$ to obtain the synthetic blurry image $\hat{\mathbf{B}}_{t}$, which is used to calculate the reconstruction loss $\mathcal{L}_{rec}$ with the given blurry frame $\mathbf{B}_{t}$, i.e.,
$\mathcal{L}_{rec}=(1-\beta)\mathcal{L}_{1}(\hat{\mathbf{B}}_{t},\mathbf{B}_{t})+\beta\mathcal{L}_{ssim}(\hat{\mathbf{B}}_{t},\mathbf{B}_{t}).$ (4)
$\mathcal{L}_{1}$ and $\mathcal{L}_{ssim}$ are $\ell_{1}$ loss and SSIM (Wang et al. 2004) loss, respectively. $\beta$ is set to $0.2$. The setting all follows 3DGS (Kerbl et al. 2023).
Continuous Camera Poses Estimation
To estimate continuous camera poses, recent methods (Zhao et al. 2024; Peng et al. 2024a; Chen and Liu 2024; Oh et al. 2024) directly optimize exposure start and end poses (i.e., $\mathbf{P}_{t,1}$ and $\mathbf{P}_{t,N}$). Then, the linear interpolation is performed between $\mathbf{P}_{t,1}$ and $\mathbf{P}_{t,N}$ to obtain the camera pose at $i$-th intermediate timestamp (i.e., $\mathbf{P}_{t,i}$),i.e.,
$\mathbf{P}_{t,i}=\mathbf{P}_{t,1}\odot\texttt{exp}(\frac{i-1}{N-1}\odot\texttt{log}(\frac{\mathbf{P}_{t,N}}{\mathbf{P}_{t,1}})).$ (5)
exp and log are exponential and logarithmic functions, respectively. $\odot$ is a pixel-wise multiply operation.
We follow the manner but deploy a tiny MLP as the camera motion predictor (see details in the Suppl) for more stable optimization. We pre-train it and static Gaussians $\mathbf{S}$ with static reconstruction loss $\mathcal{L}_{rec}^{s}$, i.e.,
$\mathcal{L}_{rec}^{s}=(1-\beta)\mathcal{L}_{1}(\hat{\mathbf{B}}_{t}^{s},\mathbf{B}_{t}^{s})+\beta\mathcal{L}_{ssim}(\hat{\mathbf{B}}_{t}^{s},\mathbf{B}_{t}^{s}).$ (6)
$\hat{\mathbf{B}}_{t}^{s}=(1-\mathbf{M}_{t})\odot\hat{\mathbf{B}}_{t}$ and $\mathbf{B}_{t}^{s}=(1-\mathbf{M}_{t})\odot\mathbf{B}_{t}$ are the static areas of $\hat{\mathbf{B}}_{t}$ and $\mathbf{B}_{t}$, respectively.
Continuous Dynamic Gaussians Estimation
We first introduce blur-aware variable canonical Gaussians for better dynamic representation at integer timestamps. Then, we describe Gaussian deformation manner. Finally, we detail how to take learnable exposure time parameters to obtain continuous dynamic Gaussians within exposure time.
Blur-Aware Variable Canonical Gaussians. Existing 4D reconstruction methods generally select a single canonical Gaussians $\mathbf{C}$ across the entire video, which may produce results with missing details in scenes with large motion. To alleviate the issue, we suggest varying the canonical Gaussians as time progresses. In such case, the $k$-th canonical Gaussians $\mathbf{C}_{k}$ is only used for some nearby timestamps, thus reducing the difficulty of motion modeling. One way to achieve this is to uniformly divide the video into $K$ segments and select $\mathbf{C}_{k}$ for $k$-th segment. Although it improves performance, selecting the one corresponding to the sharper frame is better for blur removal. In particular, we first uniformly divide the video into $K$ segments and calculate the blur level $b_{t}$ of dynamic areas for $t$-th frame following  (Bansal et al. 2016; Ren et al. 2020), i.e.,
$b_{t}=\sum_{(u,v)\in\mathbf{M}_{t}}(\Delta\mathbf{B}_{t}(u,v)-\overline{\Delta\mathbf{B}_{t}})^{2}.$ (7)
$\mathbf{M}_{t}$ indicates dynamic areas. $\Delta\mathbf{B}_{t}$ is the image Laplacian and $\overline{\Delta\mathbf{B}_{t}}$ is its mean value. The larger $b_{t}$ is, the sharper the frame is. To make the start and end frame of the segment as sharp as possible, we look for the sharp frame among their surrounding $H$ frames and redefine them as the start and end of current segment. Finally, we select the Gaussians for the sharpest frame in each segment as its canonical ones.
Gaussian Deformation. We deform dynamic Gaussians with a set of rigid transformation matrices, following Shape-of-Motion (Wang et al. 2024). Let $\{\mathbf{x}_{c},\mathbf{r}_{c},\mathbf{s},\mathbf{o},\mathbf{c}\}$, $\{\mathbf{x}_{t},\mathbf{r}_{t},\mathbf{s},\mathbf{o},\mathbf{c}\}$, and $\{\mathbf{A}_{t},\mathbf{E}_{t}\}$ denote a Gaussian in $\mathbf{C}_{k}$, the ones in $\mathbf{D}_{t}$, and the corresponding transformation matrix, respectively. It can be written as,
$\mathbf{x}_{t}=\mathbf{A}_{t}\mathbf{x}_{c}+\mathbf{E}_{t},\qquad\mathbf{r}_{t}=\mathbf{A}_{t}\mathbf{r}_{c}.$ (8)
Interpolation with Exposure Time Parameters. To get continuous dynamic Gaussians $\{\mathbf{D}_{t,i}\}_{i=1}^{N}$, one straightforward way is to deploy a series of learnable Gaussian or deformation parameters, but it is unstable to optimize. With the explicit object motion representation in eq. 8, $\mathbf{D}_{t,i}$ can be calculated by interpolating between the ones at the nearest integer timestamps, i.e.,
$\begin{split}\mathbf{D}_{t,i}=\mathbf{w}_{t,i}\odot\mathbf{D}_{t-1}+(1-\mathbf{w}_{t,i})\odot\mathbf{D}_{t},\text{ }{i}\in[1,N/2],\\ \mathbf{D}_{t,i}=(1-\mathbf{w}_{t,i})\odot\mathbf{D}_{t}+\mathbf{w}_{t,i}\odot\mathbf{D}_{t+1},\text{ }{i}\in[N/2,N].\end{split}$ (9)
$\mathbf{w}_{t,i}$ is the normalized time interval between $\mathbf{D}_{t,i}$ and $\mathbf{D}_{t,N/2}$. Thus, the problem is transformed to estimate $\mathbf{w}_{t,i}$. In the implementation, we can estimate the one at exposure start and end (i.e., $\mathbf{w}_{t,1}$ and $\mathbf{w}_{t,N}$) and then interpolate between them to get the $i$-th intermediate one $\mathbf{w}_{t,i}$, i.e.,
$\mathbf{w}_{t,i}=(1-\frac{i-1}{N-1})\odot\mathbf{w}_{t,1}+\frac{i-1}{N-1}\odot\mathbf{w}_{t,N}.$ (10)
As the object motion within the exposure can be regarded as uniform, the absolute value of $\mathbf{w}_{t,1}$ and $\mathbf{w}_{t,N}$ are equal, which is half the exposure time $\mathbf{w}_{t}$. Thus, eq. 10 can be re-written as,
$\mathbf{w}_{t,i}=(1-\frac{i-1}{N-1})\odot\frac{\mathbf{w}_{t}}{2}+\frac{i-1}{N-1}\odot(-\frac{\mathbf{w}_{t}}{2}).$ (11)
Finally, we set learnable parameters $\mathbf{w}_{t}$ for continuous dynamic Gaussians estimation within the exposure time. The canonical Gaussians, Gaussian deformation modules, and $\mathbf{w}_{t}$ are jointly optimized. The reconstruction loss for dynamic areas is similar to eq. 6.
Regularization Terms
After optimization with eq. 4, static areas of $\hat{\mathbf{I}}_{t,i}$ are sharp while dynamic areas can with notable artifacts. The reasons are below. (1) Note that multiple solutions exist for the model to fulfill eq. 4. The most ideal one is that every $\hat{\mathbf{I}}_{t,i}$ is sharp, and the most trivial one is that every $\hat{\mathbf{I}}_{t,i}$ is as blurry as $\mathbf{B}_{t}$. (2) As static areas are consistent across the entire video, the model tends to learn the underlying sharp representation for inter-frame consistency. In other words, the inter-frame consistency implicitly regularizes model optimization. To further validate this, we conduct an experiment that removes the inter-frame consistency by reducing the number of frames to one. In such a case, the static areas are blurry after optimization with eq. 4, which supports our confirmation. (3) Compared to static areas, the inter-frame consistency in dynamic ones is weaker due to object motion. It may provide insufficient regularization to guide sharp representation learning, thus leading to artifacts. To avoid this, we introduce regularization terms $\mathcal{L}_{reg}$, including exposure regularization $\mathcal{L}_{e}$, multi-frame consistency term $\mathcal{L}_{mfc}$, and multi-resolution consistency term $\mathcal{L}_{mrc}$.
First, the continuous dynamic Gaussians $\{\mathbf{D}_{t,i}\}_{i=1}^{N}$ should not be the same. In other words, the value of exposure time parameters $\mathbf{w}_{t}$ should not be too small. If $\mathbf{w}_{t}$ is too small, $\mathbf{D}_{t,i}$ is nearly the same as $\mathbf{D}_{t}$, leading to trivial solutions. We constrain $\mathbf{w}_{t}$ by $\mathcal{L}_{e}$, as,
$\mathcal{L}_{e}=\texttt{max}(0,\mathbf{\epsilon}-\mathbf{w}_{t}).$ (12)
max is the maximum function and $\mathbf{\epsilon}$ is a threshold.
Second, despite different motions, the content of multiple frames within exposure time should be similar. We utilize $\mathcal{L}_{mfc}$ to constrain consistency between neighbor frames, and that between each frame and the first frame, i.e.,
$\begin{split}\mathcal{L}_{mfc}=\frac{1}{N-1}&\sum_{i=2}^{N}(\left\|\mathbf{M}_{t,i}\odot(\hat{\mathbf{I}}_{t,{i-1\rightarrow i}}-\hat{\mathbf{I}}_{t,i})\right\|_{1}\\ &\;\;\>+\left\|\mathbf{M}_{t,1}\odot(\hat{\mathbf{I}}_{t,{i\rightarrow 1}}-\hat{\mathbf{I}}_{t,1})\right\|_{1}).\end{split}$ (13)
$\hat{\mathbf{I}}_{t,{i-1\rightarrow i}}$ and $\hat{\mathbf{I}}_{t,{i\rightarrow 1}}$ are obtained by aligning $\hat{\mathbf{I}}_{t,{i-1}}$ to $\hat{\mathbf{I}}_{t,i}$ and aligning $\hat{\mathbf{I}}_{t,i}$ to $\hat{\mathbf{I}}_{t,1}$ with a pre-trained optical flow network (Sun et al. 2018), respectively. $\mathbf{M}_{t,i}$ and $\mathbf{M}_{t,1}$ are dynamic masks for $\hat{\mathbf{I}}_{t,i}$ and $\hat{\mathbf{I}}_{t,1}$, respectively.
Third, the blur in the lower resolution is lower level and is easier to remove (Kim et al. 2022; Tao et al. 2018), thus the artifacts are less in models trained with down-sampled blurry video. Taking this advantage, we impose $\mathcal{L}_{mrc}$ to assist the optimization of high-resolution models with results from low-resolution models,i.e.,
[…]
$\hat{\mathbf{I}}_{t,i}^{l}$ is the rendered sharp image from the low-resolution model, which is pre-trained by taking the down-sampled video as supervision. $(\cdot)_{\downarrow}$ is an image down-sampling operation. sg is the stop-gradient operation.
Overall, regularization terms $\mathcal{L}_{reg}$ can be denoted as,
$\mathcal{L}_{reg}=\lambda_{e}\mathcal{L}_{e}+\lambda_{mfc}\mathcal{L}_{mfc}+\lambda_{mrc}\mathcal{L}_{mrc}.$ (15)
$\lambda_{e}$, $\lambda_{mfc}$, and $\lambda_{mrc}$ are set to 0.1, 2, 1, respectively. Besides, following Shape-of-Motion (Wang et al. 2024), we also use some other regularization terms $\mathcal{L}_{oth}$ to help reconstruct 3D motion better, and the details are in the Suppl.
Application to Multiple Tasks
The blurry videos suffer from not only motion blur, but also low frame rates and scene shake generally. Beyond novel-view synthesis, Deblur4DGS can adjust the camera poses and timestamps to address these problems, achieving video deblurring, frame interpolation, and video stabilization. First, when inputting camera poses of the blurry video, Deblur4DGS can render corresponding deblurring results. Second, when feeding the interpolated camera poses and timestamps, Deblur4DGS can produce frame-interpolated results. Third, Deblur4DGS can render a more stable video with the smoothed camera poses as inputs.
Methods
[…]
$288\times 512$
[…]
$720\times 1080$
DeformableGS 15.73 / 0.623 / 0.382 15.55 / 0.667 / 0.421
4DGaussians 21.98 / 0.801 / 0.197 21.69 / 0.831 / 0.264
E-D3DGS 23.09 / 0.830 / 0.175 22.46 / 0.844 / 0.258
Shape-of-Motion 26.06 / 0.910 / 0.144 25.81 / 0.897 / 0.246
SplineGS 26.05 / 0.901 / 0.158 24.92 / 0.883 / 0.252
DyBluRF 26.04 / 0.916 / 0.090 25.71 / 0.908 / 0.159
BARD-GS 26.91 / 0.923 / 0.077 26.34 / 0.909 / 0.139
Deblur4DGS (Ours) 27.66 / 0.935 / 0.060 27.16 / 0.927 / 0.123
Table 1: novel view synthesis results on synthetic videos.
Methods
[…]
Redmi
[…]
BARD-GS
DeformableGS 0.238 / 25.903 15.63 / 0.781 / 0.361
4DGaussians 0.236 / 26.514 21.32 / 0.863 / 0.221
E-D3DGS 0.257 / 24.967 22.69 / 0.883 / 0.217
Shape-of-Motion 0.277 / 24.538 20.53 / 0.854 / 0.289
SplineGS 0.252 / 32.022 23.93 / 0.899 / 0.197
DyBluRF 0.263 / 34.006 22.71 / 0.878 / 0.192
BARD-GS 0.288 / 35.505 22.69 / 0.874 / 0.177
Deblur4DGS (Ours) 0.356 / 36.756 23.10 / 0.879 / 0.161
Table 2: novel view synthesis results on real-world videos.
[…]
Figure 2: Visual comparisons of novel-view synthesis on real-world videos. Our method produces more photo-realistic details and less visual artifacts in both static and dynamic areas, as marked with yellow and red boxes respectively.
Experiments
Experimental Settings
Training Details. For stable optimization, we pre-train the camera motion predictor and static Gaussians $\mathbf{S}$ for $400$ epochs. After that, we jointly optimize the camera motion predictor, $\mathbf{S}$, canonical dynamic Gaussians $\{\mathbf{C}_{k}\}_{k=1}^{K}$, deformable operation $\mathcal{F}$ and exposure time parameters $\{\mathbf{w}_{t}\}_{t=1}^{T}$ for $200$ epochs. The learning rate for camera motion predictor is set to $5\times 10^{-4}$ and decayed to $1\times 10^{-5}$. The learning rate for $\{\mathbf{w}_{t}\}_{t=1}^{T}$ is set to $1\times 10^{-1}$ and decayed to $1\times 10^{-5}$. The learning rate for $\mathbf{S}$, $\{\mathbf{C}_{k}\}_{k=1}^{K}$ and $\mathcal{F}$ follows Shape-of-Motion (Wang et al. 2024). $N$ is set to 11. $K$ and $H$ are set to 5 and 3 respectively. $\mathbf{\epsilon}$ is set to $1.0$. Experiments are conducted with PyTorch (Paszke et al. 2019) on one Nvidia GeForce RTX A6000 GPU.
Evaluation Configurations. The synthetic data contains 9 scenes with significant motion blur from Stereo Blur Dataset (Zhou et al. 2019), where each scene contains blurry stereo videos and the corresponding sharp ones. Note that DyBluRF (Sun et al. 2024a) conducts experiments on $\times 2.5$ down-sampled data. We evaluate on both $\times 2.5$ down-sampled ones (i.e., $288\times 512$) and the original ones (i.e., $720\times 1080$). For novel-view synthesis and deblurring, the rendering results may be spatially misaligned with ground truth due to the calibration error of camera parameters. Thus, we first freeze the pre-trained 4D model and optimize camera poses by minimizing the photometric error between rendering results and ground truth, and then calculate metrics (i.e., PSNR, SSIM and LPIPS (Zhang et al. 2018)), following COLMAP-Free 3DGS (Fu et al. 2024). As there is no ground truth for frame interpolation and video stabilization, we employ recent no-reference metrics, i.e., CLIPIQA (Wang et al. 2023a) and MUSIQ (Ke and et al. 2021). Besides, we evaluate on 6 real-world blurry videos (i.e., Redmi data) captured by a Redmi K50 Ultra smartphone and 12 real-world ones (i.e., BARD-GS data) from BARD-GS (Lu et al. 2025), where each one contains 24 frames. For novel-view synthesis, we employ no-reference metrics (i.e., CLIPIQA and MUSIQ) and full-reference metrics (i.e., PSNR, SSIM and LPIPS) for the two data, respectively. For the other three tasks, we use no-reference metrics due to no ground truth.
Comparison with State-of-the-Art Methods
We compare with 7 state-of-the-art methods (i.e., DeformableGS (Yang et al. 2024), 4DGaussians (Wu et al. 2024), E-D3DGS (Bae et al. 2024), Shape-of-Motion (Wang et al. 2024), SplineGS (Park and et al. 2025), DyBluRF (Sun et al. 2024a) and BARD-GS (Lu et al. 2025)), where DyBluRF and BARD-GS are designed for 4D reconstruction from blurry monocular video based on NeRF and 3DGS respectively.
4D
Methods
Deblurring
Methods
[…]
$288\times 512$
[…]
$720\times 1080$
E-D3DGS None 23.09 / 0.830 / 0.175 22.46 / 0.844 / 0.258
Restormer 23.79 / 0.850 / 0.142 22.87 / 0.864 / 0.198
DSTNet 23.59 / 0.861 / 0.132 22.93 / 0.863 / 0.188
BSSTNet 23.68 / 0.855 / 0.128 23.20 / 0.877 / 0.176
Shape-of-Motion None 26.06 / 0.910 / 0.144 25.81 / 0.897 / 0.246
Restormer 26.80 / 0.917 / 0.085 26.20 / 0.911 / 0.169
DSTNet 26.60 / 0.915 / 0.082 26.08 / 0.914 / 0.140
BSSTNet 26.78 / 0.916 / 0.080 26.06 / 0.916 / 0.125
Deblur4DGS None 27.66 / 0.935 / 0.060 27.16 / 0.927 / 0.123
Table 3: Results that pre-process the blurry video with an image or video deblurring method before 4D reconstruction. ‘None’ denotes no deblurring method being used.
Novel-view synthesis. table 1 and table 2 summarize the results. First, methods (i.e., DyBluRF and BARD-GS) that perform 4D reconstruction and motion blur modeling jointly yield overall better performance, especially in LPIPS score. Although SplineGS gets better PSNR and SSIM scores in BARD-GS data, it produces blurry outputs. It is consistent with the finding in BARD-GS (Wang et al. 2024) that PSNR can sometimes yield higher values even when images appear blurrier. Second, benefiting the explicit 3D representation manner, BARD-GS outperforms DyBluRF, being the most competitive method. Third, compared with BARD-GS, our Deblur4DGS performs better due to the introduction of a series of regularization terms to avoid trivial solutions and blur-aware variable canonical Gaussians to better represent dynamic objects. Visual results in fig. 11 and fig. 2 shows that Deblur4DGS removes blur more clearly and produces less visual artifacts in both static and dynamic areas. Per-scene results and more visual results are in the Suppl.
In addition, to further demonstrate the effectiveness of Deblur4DGS, we first pre-process the blurry videos with state-of-the-art image (i.e., Restormer (Zamir et al. 2022)) or video (DSTNet (Pan et al. 2023) and BSSTNet (Zhang and et al. 2024)) deblurring methods and then perform 4D reconstruction, as shown in table 3. Compared with reconstruction from blurry videos, the incorporation of deblurring models improves performance. This is because the deblurring models remove some blur, facilitating sharp scene reconstruction. However, as the deblurring methods cannot perceive 3D structure and maintain scene geometric consistency, the reconstruction results are still unsatisfactory. In contrast, Deblur4DGS jointly reconstructs scene geometry and processes motion blur in 3D space, achieving better scene reconstruction results. Visual results are in the Suppl.
Deblurring. Apart from 4D reconstruction-based methods, we compare with some state-of-the-art image and video deblurring ones. The results are in the Suppl. Deblur4DGS obtains better results than 4D reconstruction-based methods and comparable ones to deblurring-specific ones. Compared with the former, Deblur4DGS better reconstructs the scene, thus performing better. Note that the latter ones are trained on large paired data in supervised manner while Deblur4DGS is optimized with the given blurry video in a self-supervised manner. Although the data prior makes them perform better, Deblur4DGS is more convenient to use.
Frame Interpolation. We interpolate camera poses and timestamps to generate $\times 16$ frame interpolation results. We compare with 4D reconstruction-based methods and some video frame interpolation ones (i.e., RIFE (Huang et al. 2022), EMAVFI (Zhang et al. 2023a), and VIDUE (Shang et al. 2023)). The results are in the Suppl. VIDUE is trained with large paired data in a supervised manner for joint deblurring and fame interpolation, thus achieving better results.
Video Stabilization. We employ a Gaussian filter to smooth camera poses for video stabilization, following  (Peng et al. 2024b). The results are in the Suppl. Deblur4DGS achieves pleasant scores compared with 2D video stabilization methods (i.e., MeshFlow (Liu et al. 2016) and NNDVS (Zhang et al. 2023b)) and 4D reconstruction-based ones, which benefits from better geometry reconstruction.
Figure 3: Effect of continuous camera pose (ECP) and dynamic Gaussian (EDG) estimation.
Ablation Study
We conduct experiments to validate the effectiveness of each strategy. As our strategies in Deblur4DGS are mainly designed to process dynamic areas, we exclude the pixels of static areas to calculate the metrics in dynamic ones.
Effect of ECP and EDG
ECP and EDG are introduced to process camera motion blur and object motion blur, respectively. Quantitative results and visual comparisons are shown in table 4 and fig. 3, respectively. First, without ECP and EDG, the results are almost as blurry as the input frame, as shown in fig. 3(b). Second, only with ECP, the static areas are sharp but may lead to visual artifacts in dynamic areas, as shown in fig. 3(c). It is because ECP cannot simulate the object movement. Third, we further introduce EDG to simulate that, producing visually pleasant results in both areas, as shown in fig. 3(d).
Effect of Regularization Terms
The effect of exposure regularization $\mathcal{L}_{e}$, multi-frame consistency term $\mathcal{L}_{mfc}$ and multi-resolution consistency term $\mathcal{L}_{mrc}$ are in table 5. Visual results are in the Sec.D of Suppl. Without these regularization terms, noticeable artifacts appear in dynamic regions, resulting in degraded performance. By regularizing the object motion within the exposure time distinguished, $\mathcal{L}_{e}$ improves performance. Besides, $\mathcal{L}_{mfc}$ and $\mathcal{L}_{mrc}$ additionally regularize multi-frame and multi-resolution consistency respectively, helping to alleviate artifacts. Their combinations perform best.
Effect of BAV Canonical Gaussians.
The results are summarized in table 6. First, selecting a single canonical Gaussians across the entire video (i.e., None) leads to poor performance, due to the challenge of modeling large object motion. Second, selecting variable canonical Gaussians uniformly (i.e., w/o Blur-Aware) alleviates this, leading to performance gain. We also experiment with an optical flow-based strategy (Shaw et al. ) to select canonical Gaussians, performs similar to the uniform selection. This may be due to the inaccurate estimation of optical flow from blurry images. Third, our blur-aware selection is better, as the canonical Gaussians from the sharper frame help blur removal. Visual results are in the Suppl.
ECP EDG
[…]
$288\times 512$
[…]
$720\times 1080$
$\times$ $\times$ 22.10 / 0.988 / 0.018 22.34 / 0.988 / 0.016
$\checkmark$ $\times$ 22.30 / 0.989 / 0.016 22.39 / 0.988 / 0.015
$\checkmark$ $\checkmark$ 22.36 / 0.989 / 0.015 22.63 / 0.990 / 0.014
Table 4: Effect about the estimation of continuous camera poses (ECP) and dynamic Gaussians (EDG).
$\mathcal{L}_{e}$ $\mathcal{L}_{mfc}$ $\mathcal{L}_{mrc}$
[…]
$288\times 512$
[…]
$720\times 1080$
$\times$ $\times$ $\times$ 21.87 / 0.987 / 0.017 22.30 / 0.988 / 0.016
$\times$ $\checkmark$ $\checkmark$ 22.30 / 0.989 / 0.015 22.56 / 0.989 / 0.015
$\checkmark$ $\times$ $\times$ 22.01 / 0.988 / 0.015 22.40 / 0.989 / 0.015
$\checkmark$ $\checkmark$ $\times$ 22.16 / 0.989 / 0.016 22.49 / 0.989 / 0.015
$\checkmark$ $\times$ $\checkmark$ 22.22 / 0.989 / 0.015 22.54 / 0.989 / 0.014
$\checkmark$ $\checkmark$ $\checkmark$ 22.36 / 0.989 / 0.015 22.63 / 0.990 / 0.014
Table 5: Effetct of regularization terms (see eq. 15).
Methods
[…]
$288\times 512$
[…]
$720\times 1080$
None 22.13 / 0.988 / 0.017 22.30 / 0.988 / 0.016
w/o Blur-Aware 22.29 / 0.989 / 0.016 22.57 / 0.989 / 0.015
Ours 22.36 / 0.989 / 0.015 22.63 / 0.990 / 0.014
Table 6: Effect of blur-aware variable (BAV) canonical Gaussians. ‘None’ denotes selecting a single one.
Conclusions
In this work, we propose Deblur4DGS, a 4D Gaussian Splatting framework to reconstruct a high-quality 4D model from blurry monocular video. In particular, with the explicit motion trajectory modeling, we propose to transform the challenging continuous dynamic representation estimation within an exposure time into the exposure time estimation, where a series of regularizations are suggested to tackle the under-constrained optimization. Besides, a blur-aware variable canonical Gaussians is present to represent objects with large motion better. Beyond novel-view synthesis, Deblur4DGS can improve blurry video quality from multiple perspectives, including deblurring, frame interpolation, and video stabilization. Extensive results show Deblur4DGS outperforms state-of-the-art 4D reconstruction methods.
Acknowledgements
This work was supported by the National Natural Science Foundation of China under Grant No. 62371164 and the National Key RD Program of China under Grant No. 2022YFA1004100.
Supplementary Material
The content of the supplementary material involves:
  • Structure of camera motion predictor in Sec.A.
  • Other regularization terms in Sec.B.
  • More comparison results in Sec.C.
  • More ablation results in Sec.D.
A Structure of Camera Motion Predictor
The structure of camera motion predictor is provided in fig. 4. It first embeds the camera pose $\mathbf{P}_{t}$ to a higher dimensional space using high frequency encoder (Mildenhall et al. 2021). Then, three FC blocks are stacked, each consisting of an FC layer followed by a ReLU operation. Finally, we deploy two heads to predict the camera pose at exposure start and end (i.e., $\mathbf{P}_{t,1}$ and $\mathbf{P}_{t,N}$), respectively.
B Other Regularization Terms
Following Shape-of-Motion (Wang et al. 2024), we use some other regularization terms $\mathcal{L}_{oth}$ to help reconstruct 3D motion better, including mask regularization $\mathcal{L}_{mask}$, 2D tracks regularization $\mathcal{L}_{track}$ and distance-preserving regularization $\mathcal{L}_{rigid}$.
Specifically, we render the masks within the exposure time $\{\hat{\mathbf{M}}_{t,i}\}_{i=1}^{N}$ to indicate dynamic areas. To supervise the training, we synthesize the mask $\hat{\mathbf{M}}_{t}^{B}$ for the synthetic blurry image $\hat{\mathbf{B}}_{t}$ as,
$\hat{\mathbf{M}}_{t}^{B}(u,v)=\texttt{max}\{\hat{\mathbf{M}}_{t,1}(u,v),\hat{\mathbf{M}}_{t,2}(u,v),...,\hat{\mathbf{M}}_{t,N}(u,v)\}.$ (16)
$(u,v)$ is the pixel location. The mask regularization $\mathcal{L}_{mask}$ can be written as,
$\mathcal{L}_{mask}=\mathcal{L}_{1}(\hat{\mathbf{M}}_{t}^{B},\mathbf{M}_{t}),$ (17)
where $\mathbf{M}_{t}$ is the mask obtained by applying SAM2 (Ravi et al. 2024) to the ground truth blurry frame. Besides, we render the 2D tracks $\hat{\mathbf{U}}_{t\rightarrow t^{\prime}}$ from a pair of randomly sampled query time $t$ and target time $t^{\prime}$. We supervise it by the lifted long-range 2D tracks $\mathbf{U}_{t\rightarrow t^{\prime}}$ that are extracted from TAPIR (Doersch and et al. 2023), i.e.,
$\mathcal{L}_{track}=\mathcal{L}_{1}(\hat{\mathbf{U}}_{t\rightarrow t^{\prime}},\mathbf{U}_{t\rightarrow t^{\prime}}).$ (18)
Finally, we enforce a distance-preserving loss $\mathcal{L}_{rigid}$ between randomly sampled dynamic Gaussians and their $J$-nearest neighbors. Let $\mathbf{x}_{t}$ and $\mathbf{x}_{t^{\prime}}$ denote the position of a Gaussian at time $t$ and $t^{\prime}$, and $\mathcal{C}_{J}(\mathbf{x}_{t})$ denote the set of $J$-nearest neighbors of $\mathbf{x}_{t}$. $\mathcal{L}_{rigid}$ can be written as,
[…]
dist measures the Euclidean distance.
Figure 4: Structure of camera motion predictor.
Overall, $\mathcal{L}_{oth}$ can be written as,
$\mathcal{L}_{oth}=\lambda_{mask}\mathcal{L}_{mask}+\lambda_{track}\mathcal{L}_{track}+\lambda_{rigid}\mathcal{L}_{rigid}.$ (20)
$\lambda_{mask}$, $\lambda_{track}$ and $\lambda_{rigid}$ are set to 1, 2, and 2, respectively.
Figure 5: Effect of regularization terms $\mathcal{L}_{reg}$. $\mathcal{L}_{reg}$ includes exposure regularization $\mathcal{L}_{e}$, multi-frame consistency regularization $\mathcal{L}_{mfc}$, and multi-resolution consistency regularization $\mathcal{L}_{mrc}$.
Figure 6: Some examples of our captured real-world blurry videos by the Redmi K50 Ultra smartphone.
[…]
Figure 7: Effect of BAV canonical Gaussians.
[…]
Figure 8: Visual comparisons with methods that pre-process the blurry video with an image (i.e., Restormer (Zamir et al. 2022)) or video (DSTNet (Pan et al. 2023) and BSSTNet (Zhang and et al. 2024)) deblurring method before performing 4D reconstruction (i.e., Shape-of-Motion (Wang et al. 2024)).
[…]
Figure 9: Example of masks in extreme blurry videos. SAM 2 (Ravi and et al. 2024) may fail to capture small dynamic objects under such extreme blur, as shown in the red box.
C More Comparison Results
fig. 6 shows some examples of real-world videos captured by the Redmi K50 Ultra smartphone. We evaluate methods on four tasks, i.e., novel-view synthesis, deblurring, frame interpolation, and video stabilization. Per-scene results for novel-view synthesis are summarized in table 7, table 8, table 9 and table 10. Rendering speed comparisons are in table 11. More visual comparisons for novel-view synthesis are in fig. 10 and fig. 11. They show that our Deblur4DGS produces more photo-realistic details and fewer visual artifacts. Additionally, the visual comparisons with methods that pre-process the blurry video with an image (i.e., Restormer (Zamir et al. 2022)) or video (DSTNet (Pan et al. 2023) and BSSTNet (Zhang and et al. 2024)) deblurring method before 4D reconstruction (i.e., Shape-of-Motion (Wang et al. 2024)) are in fig. 8, which shows that our Deblur4DGS still performs better.
Besides, table 12, table 14 and table 16 summarize the results for deblurring, frame interpolation, and video stabilization on synthetic videos. table 13, table 15, and table 17 summarize those on real-world videos. Deblur4DGS outperforms state-of-the-art 4D reconstruction methods and has competitive capabilities in comparison with task-specific video processing models trained in a supervised manner. Furthermore, visual results for deblurring in fig. 12 show that compared with 4D reconstruction methods, Deblur4DGS produces sharper contents. Moreover, the visual results of images within an exposure time and the corresponding synthetic blurry image are in fig. 13. It shows that Deblur4DGS successfully synthesize the blur by estimating camera motion and object motion trajectory. We provide some videos at https://deblur4dgs.github.io/.
D More Ablation Results
Visual results in fig. 5 validate the effectiveness of the proposed regularization terms. Without these regularization terms, notable artifacts appear in dynamic areas. By regularizing the object motion within the exposure time distinguished, $\mathcal{L}_{e}$ improves performance. Besides, $\mathcal{L}_{mfc}$ and $\mathcal{L}_{mrc}$ additionally regularize multi-frame and multi-resolution consistency respectively, helping to alleviate artifacts. In addition, visual results in fig. 7 validate the effectiveness of blur-aware variable (BAV) canonical Gaussians. Note that we extract dynamic masks using SAM 2 (Ravi and et al. 2024), which is generally robust enough to produce accurate masks even in blurry videos. However, under extreme blur, it may fail to capture small dynamic objects, as shown in fig. 9.
Table 7: Per-Scene results for novel view synthesis on $288\times 512$ synthetic data.
Methods
[…]
Skating
[…]
Seesaw
[…]
Street
[…]
Basketball
[…]
Children
[…]
Sailor
[…]
Third
[…]
Man
[…]
Women
DeformableGS 16.74/0.557/0.482 18.27/0.695/0.237 12.13/0.529/0.587 15.34/0.645/0.306 15.04/0.598/0.472 12.89/0.491/0.539 16.91/0.725/0.284 19.48/0.747/0.273 14.66/0.618/0.257
4DGaussians 24.31/0.870/0.130 24.42/0.887/0.104 17.58/0.605/0.326 25.32/0.897/0.117 18.96/0.691/0.344 18.98/0.722/0.323 27.47/0.937/0.089 24.49/0.872/0.136 16.32/0.735/0.209
E-D3DGS 25.59/0.893/0.111 25.16/0.893/0.111 18.37/0.677/0.275 24.71/0.882/0.129 22.34/0.809/0.255 19.91/0.747/0.300 28.56/0.939/0.087 24.52/0.873/0.137 18.67/0.755/0.185
Shape-of-Motion 30.95/0.959/0.088 26.91/0.929/0.086 28.56/0.947/0.073 23.31/0.865/0.180 23.31/0.841/0.306 22.07/0.894/0.190 29.89/0.960/0.074 25.82/0.947/0.147 23.72/0.851/0.151
SplineGS 28.55/0.948/0.095 28.48/0.953/0.075 24.55/0.924/0.082 26.02/0.915/0.129 22.21/0.777/0.438 25.31/0.876/0.232 29.65/0.960/0.076 25.89/0.908/0.149 23.83/0.848/0.129
DyBluRF 29.54/0.935/0.076 26.46/0.936/0.080 26.40/0.938/0.085 26.24/0.907/0.059 25.32/0.903/0.111 25.27/0.905/0.123 28.57/0.937/0.070 24.40/0.926/0.103 22.23/0.860/0.105
BARD-GS 29.78/0.940/0.060 27.12/0.944/0.057 28.25/0.939/0.067 26.25/0.923/0.083 25.67/0.910/0.109 25.21/0.901/0.111 29.56/0.961/0.051 26.87/0.934/0.067 23.56/0.858/0.087
Deblur4DGS (Ours) 30.84/0.965/0.047 28.45/0.954/0.049 29.01/0.953/0.048 27.00/0.932/0.065 26.05/0.916/0.087 25.72/0.910/0.097 29.82/0.964/0.047 28.53/0.957/0.032 23.55/0.865/0.066
Table 8: Per-Scene results for novel view synthesis on $720\times 1280$ synthetic data.
Methods
[…]
Skating
[…]
Seesaw
[…]
Street
[…]
Basketball
[…]
Children
[…]
Sailor
[…]
Third
[…]
Man
[…]
Women
DeformableGS 16.58/0.708/0.482 17.36/0.739/0.332 11.53/0.605/0.673 15.34/0.730/0.324 16.15/0.744/0.587 12.36/0.612/0.588 17.46/0.786/0.345 19.50/0.782/0.323 14.22/0.646/0.369
4DGaussians 23.75/0.880/0.203 23.84/0.886/0.190 17.99/0.740/0.344 20.66/0.827/0.193 22.32/0.830/0.398 18.89/0.784/0.382 26.99/0.930/0.168 23.62/0.852/0.210 17.16/0.752/0.286
E-D3DGS 25.09/0.897/0.186 24.32/0.891/0.183 18.04/0.764/0.353 23.87/0.873/0.172 22.27/0.833/0.394 19.05/0.791/0.376 27.37/0.927/0.181 23.71/0.853/0.208 18.43/0.771/0.276
Shape-of-Motion 30.11/0.946/0.169 25.57/0.913/0.200 27.87/0.939/0.139 23.21/0.876/0.224 24.12/0.860/0.488 23.76/0.864/0.355 28.92/0.948/0.156 24.84/0.879/0.238 23.88/0.853/0.248
SplineGS 28.30/0.938/0.176 27.27/0.936/0.174 24.07/0.919/0.146 24.79/0.895/0.199 22.24/0.790/0.440 20.93/0.816/0.472 28.79/0.948/0.155 24.76/0.875/0.240 23.06/0.833/0.268
DyBluRF 29.14/0.915/0.134 26.56/0.930/0.128 26.01/0.928/0.135 25.84/0.902/0.147 25.12/0.893/0.245 24.97/0.903/0.220 27.17/0.928/0.124 24.30/0.919/0.151 22.23/0.851/0.153
BARD-GS 29.30/0.930/0.122 25.89/0.935/0.125 27.98/0.942/0.122 25.96/0.912/0.116 24.29/0.831/0.250 24.70/0.880/0.193 28.90/0.948/0.112 25.84/0.920/0.101 24.20/0.890/0.113
Deblur4DGS (Ours) 30.77/0.956/0.093 27.52/0.942/0.107 28.22/0.944/0.093 26.93/0.931/0.109 25.68/0.905/0.235 24.91/0.897/0.185 28.96/0.955/0.104 28.05/0.948/0.064 23.50/0.865/0.114
Table 9: Per-Scene results for novel view synthesis on real-world videos captured by Redmi K50 Ultra.
Methods
[…]
Girl
[…]
Running
[…]
Boy
[…]
Walking
[…]
Airport
[…]
Bookshop
DeformableGS 0.204/26.941 0.190/22.321 0.266/25.099 0.246/30.193 0.301/23.545 0.219/27.321
4DGaussians 0.220/31.099 0.201/26.140 0.270/21.963 0.237/27.717 0.267/22.642 0.224/29.499
E-D3DGS 0.232/30.962 0.219/25.331 0.287/23.929 0.257/30.423 0.282/23.707 0.267/33.447
Shape-of-Motion 0.267/31.979 0.234/22.515 0.322/20.214 0.241/27.609 0.344/18.093 0.252/26.827
SplineGS 0.228/48.300 0.207/34.833 0.308/29.323 0.248/26.299 0.299/25.081 0.222/28.295
DyBluRF 0.169/21.756 0.160/29.160 0.303/31.051 0.321/25.824 0.218/38.048 0.257/42.090
BARD-GS 0.245/46.379 0.226/30.318 0.349/35.593 0.270/33.241 0.287/30.259 0.351/37.244
Deblur4DGS (Ours) 0.315/43.292 0.250/31.654 0.467/38.602 0.342/37.978 0.433/30.573 0.327/38.435
Table 10: Per-Scene results for novel view synthesis on real-world videos from BARD-GS (Lu et al. 2025).
Methods
[…]
card
[…]
cube-desk
[…]
kitchen
[…]
micro-lab
[…]
pen-spinning
[…]
poster
DeformableGS 13.85/0.762/0.394 13.95/0.749/0.394 16.04/0.812/0.357 17.21/0.769/0.322 15.33/0.677/0.369 21.74/0.865/0.317
4DGaussians 22.39/0.899/0.156 24.69/0.896/0.207 21.05/0.876/0.280 21.34/0.852/0.249 18.98/0.766/0.380 30.34/0.945/0.163
E-D3DGS 21.65/0.882/0.155 24.83/0.899/0.215 21.52/0.888/0.285 22.95/0.886/0.228 18.91/0.769/0.391 30.55/0.946/0.168
Shape-of-Motion 18.98/0.863/0.340 21.26/0.884/0.308 21.26/0.885/0.308 24.80/0.882/0.234 15.06/0.721/0.567 30.50/0.950/0.154
SplineGS 22.64/0.902/0.128 24.67/0.893/0.262 24.54/0.913/0.256 25.86/0.897/0.227 21.19/0.826/0.253 30.57/0.951/0.159
DyBluRF 20.73/0.841/0.169 21.41/0.825/0.179 21.78/0.849/0.191 27.04/0.863/0.116 21.60/0.652/0.345 30.15/0.925/0.088
BARD-GS 20.54/0.851/0.211 22.39/0.863/0.203 21.49/0.883/0.190 24.41/0.881/0.150 21.45/0.812/0.216 29.54/0.954/0.090
Deblur4DGS (Ours) 21.56/0.882/0.189 24.70/0.897/0.155 24.31/0.916/0.144 23.27/0.863/0.145 19.89/0.737/0.243 29.38/0.945/0.089
Methods
[…]
rubik-cube
[…]
shark-spin
[…]
tennis-ball
[…]
toycar
[…]
walk
[…]
windmill
DeformableGS 14.38/0.827/0.335 18.09/0.837/0.299 15.19/0.708/0.362 13.08/0.768/0.347 16.21/0.787/0.342 15.57/0.800/0.359
4DGaussians 20.23/0.905/0.188 23.11/0.896/0.221 21.35/0.838/0.189 19.10/0.887/0.155 25.01/0.905/0.163 20.62/0.792/0.257
E-D3DGS 20.10/0.908/0.200 23.07/0.897/0.235 22.52/0.852/0.185 20.15/0.890/0.145 24.06/0.899/0.204 21.50/0.894/0.204
Shape-of-Motion 15.10/0.847/0.354 11.37/0.724/0.468 20.63/0.830/0.182 18.32/0.878/0.168 23.91/0.893/0.222 21.70/0.890/0.231
SplineGS 20.78/0.907/0.196 25.08/0.910/0.251 20.47/0.831/0.173 22.35/0.928/0.095 25.25/0.909/0.178 23.67/0.915/0.185
DyBluRF 19.65/0.829/0.214 23.87/0.889/0.194 20.99/0.731/0.224 19.83/0.845/0.162 24.46/0.821/0.180 20.89/0.784/0.249
BARD-GS 21.09/0.900/0.169 23.90/0.901/0.252 19.16/0.745/0.196 19.31/0.878/0.191 24.51/0.895/0.138 24.48/0.927/0.123
Deblur4DGS (Ours) 21.05/0.917/0.156 25.09/0.915/0.162 19.96/0.777/0.179 18.80/0.864/0.212 25.60/0.916/0.112 23.55/0.916/0.141
Table 11: Rendering speed comparisons on $720\times 1080$ images.
Methods DeformableGS 4DGaussians E-D3DGS Shape-of-Motion SplineGS DyBluRF BARD-GS Deblur4DGS
Rendeing Speed (FPS) 98.03 56.82 81.31 96.22 130.12 0.04 80.43 96.22
Table 12: deblurring results on synthetic datasets.
[…]
288×512 720×1080
Restormer 35.45 / 0.984 / 0.023 34.66 / 0.976 / 0.037
DSTNet 34.79 / 0.981 / 0.020 33.90 / 0.973 / 0.034
BSSTNet 35.51 / 0.985 / 0.016 34.89 / 0.980 / 0.025
DeformableGS 26.88 / 0.866 / 0.177 25.53 / 0.849 / 0.239
4DGaussians 29.21 / 0.916 / 0.132 28.41 / 0.914 / 0.201
E-D3DGS 29.79 / 0.932 / 0.122 28.35 / 0.914 / 0.206
Shape-of-Motion 28.11 / 0.935 / 0.150 27.27 / 0.912 / 0.240
SplineGS 27.09 / 0.916 / 0.153 26.26 / 0.908 / 0.222
DyBluRF 29.44 / 0.947 / 0.081 28.31 / 0.916 / 0.127
BARD-GS 29.98 / 0.949 / 0.080 28.28 / 0.919 / 0.129
Deblur4DGS (Ours) 30.36 / 0.955 / 0.078 29.53 / 0.929 / 0.109
Table 13: deblurring results on real-world datasets.
Methods $uparrow CLIPIQA\uparrow/MUSIQ\uparrow$
Redmi BARD-GS
Restormer 0.254 / 37.299 0.278 / 40.378
DSTNet 0.260 / 40.213 0.217 / 27.741
BSSTNet 0.266 / 47.639 0.280 / 45.798
DeformableGS 0.249 / 26.894 0.325 / 31.186
4DGaussians 0.241 / 27.302 0.337 / 31.409
E-D3DGS 0.260 / 28.198 0.317 / 26.567
Shape-of-Motion 0.277 / 24.846 0.311 / 27.872
SplineGS 0.263 / 32.044 0.329 / 36.453
DyBluRF 0.236 / 35.240 0.295 / 34.792
BARD-GS 0.290 / 35.704 0.392 / 39.983
Deblur4DGS (Ours) 0.358 / 37.051 0.409 / 41.318
Table 14: frame interpolation results on synthetic datasets.
Methods $uparrow CLIPIQA\uparrow / MUSIQ\uparrow$
288×512 720×1080
RIFE 0.178 / 41.189 0.156 / 31.274
EMAVFI 0.179 / 39.784 0.174 / 31.031
VIDUE 0.263 / 61.333 0.186 / 49.195
DeformableGS 0.169 / 38.312 0.172 / 29.264
4DGaussians 0.171 / 41.018 0.180 / 32.032
E-D3DGS 0.173 / 39.809 0.184 / 31.145
Shape-of-Motion 0.176 / 38.871 0.202 / 31.255
SplineGS 0.184 / 44.963 0.194 / 37.788
DyBluRF 0.149 / 50.689 0.125 / 35.456
BARD-GS 0.198 / 51.123 0.196 / 38.256
Deblur4DGS (Ours) 0.201 / 52.852 0.207 / 39.721
Table 15: frame interpolation results on real-world datasets.
Methods $uparrow CLIPIQA\uparrow / MUSIQ\uparrow$
Redmi BARD-GS
RIFE 0.257 / 30.053 0.300 / 30.781
EMAVFI 0.258 / 27.025 0.318 / 30.082
VIDUE 0.273 / 40.332 0.346 / 45.782
DeformableGS 0.242 / 25.972 0.320 / 30.504
4DGaussians 0.234 / 26.923 0.333 / 31.010
E-D3DGS 0.249 / 26.425 0.314 / 25.974
Shape-of-Motion 0.286 / 28.078 0.313 / 27.809
SplineGS 0.272 / 33.553 0.330 / 37.887
DyBluRF 0.230 / 35.523 0.291 / 34.217
BARD-GS 0.301 / 36.003 0.390 / 40.002
Deblur4DGS (Ours) 0.360 / 37.224 0.416 / 41.933
Table 16: video stabilization results on synthetic datasets.
Methods $uparrow CLIPIQA\uparrow/MUSIQ\uparrow$
288×512 720×1080
MeshFlow 0.154 / 34.416 0.136 / 31.105
NNDVS 0.180 / 43.003 0.128 / 32.262
DeformableGS 0.169 / 37.978 0.173 / 28.811
4DGaussians 0.151 / 41.662 0.177 / 32.623
E-D3DGS 0.178 / 40.773 0.185 / 32.362
Shape-of-Motion 0.175 / 38.971 0.202 / 31.183
SplineGS 0.184 / 43.834 0.185 / 36.993
DyBluRF 0.148 / 32.091 0.126 / 36.115
BARD-GS 0.196 / 51.479 0.193 / 38.472
Deblur4DGS (Ours) 0.201 / 53.060 0.206 / 39.786
Table 17: video stabilization results on real-world datasets.
Methods $uparrow CLIPIQA\uparrow/MUSIQ\uparrow$
Redmi BARD-GS
MeshFlow 0.331 / 28.650 0.347 / 33.324
NNDVS 0.254 / 28.844 0.317 / 33.674
DeformableGS 0.235 / 25.309 0.318 / 30.144
4DGaussians 0.231 / 26.472 0.328 / 31.487
E-D3DGS 0.253 / 27.454 0.317 / 26.863
Shape-of-Motion 0.269 / 24.142 0.311 / 27.875
SplineGS 0.259 / 31.766 0.328 / 36.634
DyBluRF 0.260 / 34.117 0.292 / 34.319
BARD-GS 0.295 / 35.892 0.387 / 39.275
Deblur4DGS (Ours) 0.352 / 36.351 0.408 / 41.317
[…]
Figure 10: Visual comparisons of novel-view synthesis on the $720\times 1080$ images. Our method produces more photo-realistic details in both static and dynamic areas, as marked with yellow and red boxes respectively.
[…]
Figure 11: Visual comparisons of novel-view synthesis on real-world videos. Our method produces more photo-realistic details in both static and dynamic areas, as marked with yellow and red boxes respectively.
[…]
Figure 12: Visual comparisons of deblurring on the $720\times 1080$ images. Compared with 4D reconstruction-based methods, Deblur4DGS produces sharper contents and fewer artifacts in both static and dynamic areas, as marked with yellow and red boxes respectively.
[…]
Figure 13: Visual results of images within an exposure time and the synthetic blurry image. The red line is a horizontal reference line. We indicate some regions for easier observation with yellow arrows